### v1

# TimeRecipe: A Time-Series Forecasting Recipe via Benchmarking Module Level Effectiveness

Abstract
Time-series forecasting is an essential task with wide real-world applications across domains. While recent advances in deep learning have enabled time-series forecasting models with accurate predictions, there remains considerable debate over which architectures and design components, such as series decomposition or normalization, are most effective under varying conditions. Existing benchmarks primarily evaluate models at a high level, offering limited insight into why certain designs work better. To mitigate this gap, we propose TimeRecipe, a unified benchmarking framework that systematically evaluates time-series forecasting methods at the module level. TimeRecipe conducts over 10,000 experiments to assess the effectiveness of individual components across a diverse range of datasets, forecasting horizons, and task settings. Our results reveal that exhaustive exploration of the design space can yield models that outperform existing state-of-the-art methods and uncover meaningful intuitions linking specific design choices to forecasting scenarios. Furthermore, we release a practical toolkit within TimeRecipe that recommends suitable model architectures based on these empirical insights. The benchmark is available at: https://github.com/AdityaLab/TimeRecipe.
1 Introduction
Time-series forecasting plays a critical role in a wide range of real-world domains, including economics, urban computing, and epidemiology zhu2002statstream ; zheng2014urban ; deb2017review ; mathis2024evaluation ; liu2025cape . These applications focus on predicting future trends or events based on patterns observed in historical time-series data. Recently, the emergence of deep learning has significantly advanced the field, leading to the development of numerous forecasting models lai2018modeling ; torres2021deep ; salinas2020deepar ; Yuqietal-2023-PatchTST ; zhou2021informer . These models have demonstrated strong predictive accuracy and generalization capability across diverse datasets and domains, particularly within the supervised time-series forecasting paradigm.
Despite recent successes, particularly at the model level (i.e., end-to-end forecasting architectures and pipelines), ongoing debates persist regarding the most effective deep learning strategies at the module level, referring to the internal design components within forecasting models. For example, while Transformer-based architectures are known for their ability to capture long-range temporal dependencies, they tend to struggle to generalize well on highly irregular time-series patterns, such as ETT time series. This has led to the reflection in reconsidering the effectiveness of MLP-based designs zeng2023transformers ; xu2023fits ; wang2024timemixer ; ni2025timedistill . Beyond architectural debates, further divergence arises in the design of specific modules and components. For instance, wang2024timemixer++ highlights the importance of separately modeling trend and seasonal components via series decomposition, which is overlooked in other contemporary works shi2024time . Additionally, zhao2023performative raises concerns for the effectiveness of instance normalization in short-term forecasting tasks when capturing rapid trend changes, despite it being a commonly effective practice primarily associated with long-term forecasting scenarios.
However, despite ongoing debates surrounding divergent design choices at module level, existing benchmarks only emphasize evaluation at the model level qiu2024tfb ; aksu2024gift ; li2024foundts ; du2024tsi , overlooking the importance of benchmarking the effectiveness of specific module-level components. These studies typically conclude with case-specific best-performing models, which can be less practical when applying benchmarked results to real-world forecasting scenarios that fall outside the benchmarked settings brigato2025position . Moreover, while these benchmarks are comprehensive and empirically informative, they offer limited intuitive insights into why certain models outperform others in specific forecasting scenarios, leaving important questions about model behavior and design choices underexplored.
To bridge this gap and gain a deeper understanding of the underlying factors driving model performance, we propose TimeRecipe. While existing benchmarks focus primarily on model-level evaluation and offer limited interpretability, TimeRecipe takes a step further by benchmarking time-series forecasting methods at a finer-grained, module level. Specifically, TimeRecipe aims to assess the effectiveness of individual components commonly used in state-of-the-art forecasting models. This enables us to answer a key research question: Which modules and model designs are most effective under specific time-series forecasting scenarios? Guided by this question, we summarize the main contributions of this work as follows:
  • Novel Benchmarking Scope: Unlike existing time-series forecasting benchmarks that focus primarily on holistic evaluation of entire state-of-the-art (SOTA) models, TimeRecipe focuses on assessing the effectiveness of individual modules commonly used in model construction. To the best of our knowledge, TimeRecipe is the first benchmark to systematically explore the design space of forecasting models for supervised time-series forecasting tasks at the modular level.
  • Comprehensive Evaluations: TimeRecipe evaluates hundreds of module combinations across diverse forecasting scenarios, spanning univariate and multivariate settings, as well as short- and long-term horizons. The benchmark encompasses dozens of datasets and involves over 10,000 distinct experiments, offering a robust and exhaustive evaluation framework.
  • Insightful Findings: The TimeRecipe benchmark reveals that exhaustive exploration of the modular design space can yield forecasting models that outperform existing SOTA approaches. Moreover, it uncovers meaningful correlations between module effectiveness and specific characteristics of time-series data and forecasting tasks, offering insights beyond raw accuracy.
  • Actionable Toolkit: Building on the above insights, we develop a training-free toolkit that makes model architecture selections within TimeRecipe. We demonstrate its effectiveness by comparing the selected architectures against those discovered via exhaustive search.
Problem Formulation. We define the time-series forecasting problem following popular existing formulations zhou2021informer ; liu2023itransformer ; liu2024time : Given historical observations $\text{X}_{t}=\{\bm{x}_{t-L},\ldots,\bm{x}_{t}\}\in\mathbb{R}^{L\times d_{\text{X}}}$ consisting of $L$ past time steps and $d_{\text{X}}$ variables at time step $t$, the goal is to predict the future $H$ steps $\text{Y}_{t}=\{\bm{x}_{t+1},\ldots,\bm{x}_{t+H}\}\in\mathbb{R}^{H\times d_{\text{X}}}$. For convenience, we denote X as the collection of all $\text{X}_{t}$ over the full time series of length $T$, and similarly, Y as the collection of all corresponding $\text{Y}_{t}$. The time-series forecasting task aims to learn a model parameterized by $\theta$ through empirical risk minimization (ERM) to obtain $f_{\theta}:\text{X}\rightarrow\text{Y}$ for all time steps $t\in T$.
Additional Related Work. Due to page limitations, we provide an extended discussion of related works in Appendix A, including reviews of model designs for supervised learning and foundation models in time-series forecasting, as well as existing time-series forecasting benchmarks.
2 TimeRecipe Framework
Since the introduction of Transformer-based architectures to time-series forecasting, particularly following the release of Informer zhou2021informer , most modern approaches have converged toward a common design paradigm, referred to here as the Canonical Architecture, as illustrated in Figure 1. This canonical architecture consists of five major components: pre-processing, embedding, feed-forward modeling, projection, and post-processing (which is paired with pre-processing operations). The canonical architecture serves as the foundation of TimeRecipe and captures the typical structure adopted by many state-of-the-art (SOTA) forecasting methods.
Figure 1: The proposed canonical architecture in TimeRecipe for constructing general time-series forecasting models. The canonical architecture comprises five key components: pre-processing, embedding, feed-forward modeling, projection, and post-processing.
Following this structure, TimeRecipe is designed to offer a systematic evaluation with in-depth intuitions of alternative module choices across key components of the canonical architecture. Specifically, TimeRecipe benchmarks common module designs for pre-processing, embedding, and feed-forward modeling, with the aim of understanding their relative effectiveness across diverse forecasting scenarios and in conjunction with other design choices. For example, one example question to answer by our benchmark could be: how do different embedding strategies perform under varying task settings, and how do they interact with other designs from other components of the canonical architecture?
We omit the benchmarking of the projection component, as it typically consists of a simple linear layer and has attracted less interest regarding its design effectiveness. Similarly, the post-processing component is inherently paired with pre-processing operations; thus, evaluating the effectiveness of pre-processing modules simultaneously assesses the corresponding post-processing steps. We detail all benchmarked module designs included by TimeRecipe in the subsequent sections.
2.1 Pre-Processing
In TimeRecipe, we consider two types of popular time series pre-processing approaches that are widely used in time-series forecasting tasks: Instance Normalization kim2021reversible ; liu2022non ; fan2023dish ; liu2023adaptive ; han2024sin and Series Decomposition wu2021autoformer ; wang2024timemixer ; wang2024timemixer++ .
Instance Normalization. Instance normalization normalizes each input sample independently to a standard 0–1 distribution, regardless of its original distribution. This process enables the model to learn translations from historical lookback to horizon predictions more stably, as all samples are mapped into a consistent distributional space. Instance normalization is paired with a denormalization step applied after the model output, projecting the predictions from the normalized space back to the original feature space. The full process is as follows:
[…]
Series Decomposition. Series decomposition in time-series forecasting aims to disentangle the seasonal and trend components within input instances using simple moving average operations. The trend component, obtained via moving average, captures the overall directional changes of the time series and primarily consists of low-frequency variations. The seasonal component is defined as the residual between the original time series and the extracted trend, typically reflecting higher-frequency, periodic patterns. We formulate this process as follows.
$\text{X}_{t}^{\text{Trend}}=\text{AvgPool}(\text{Padding}(\text{X}_{t})),\>\>\text{X}_{t}^{\text{Season}}=\text{X}_{t}-\text{X}_{t}^{\text{Trend}}$ (2)
[…]
2.2 Embedding
In TimeRecipe, we consider four popular embedding approaches widely adopted in time-series forecasting tasks: Token zhou2021informer , Patch Yuqietal-2023-PatchTST , Invert liu2023itransformer , and Frequency xu2023fits Embedding. Additionally, we include a no-embedding variant as a controlled baseline for completeness, although it typically yields inferior performance, e.g., applying feed-forward modeling directly on the raw feature space. In short, the mathematical formulations of these embeddings are summarized in Equation 3.
$\text{X}_{t}\in[\text{B},\text{L},\text{D}],$ $\text{X}_{t}^{\text{Token}}$ $=\text{Conv}(\text{Padding}(\text{X}_{t}^{T}))^{T}$ $\in[\text{B},\text{L},\text{H}]$ (3)
$\text{X}_{t}^{\text{Patch}}$ $=\text{Conv}_{k=\text{PatchLen},s=\text{Stride}}(\text{Padding}(\text{X}_{t}^{T}))$ $\in[\text{B}\times\text{D},\text{H},\lfloor\frac{\text{L}-\text{PatchLen}}{\text{Stride}}\rfloor+2]$
$\text{X}_{t}^{\text{Invert}}$ $=\text{Linear}(\text{X}_{t}^{T})$ $\in[\text{B},\text{H},\text{D}]$
$\text{X}_{t}^{\text{Freq}}$ $=\text{rFFT}(\text{X}_{t})$ $\in[\text{B},\lfloor\frac{\text{L}}{2}\rfloor+1,\text{D}]$
Here, the superscript $T$ denotes the transpose operation between the second and third dimensions. D is the number of raw features (i.e., $d_{\text{X}}$), and H represents the hidden dimension. We detail the specifics of each embedding strategy in the following subsections.
Token Embedding. Token embedding applies convolutional operations along the temporal axis of the input time series to project each timestamp into a higher-dimensional embedding space, analogous to word embeddings in natural language processing devlin2019bert . Specifically, each timestamp, comprising all features and neighboring time steps (depending on the kernel size), is treated as an individual token. This approach preserves temporal order while enabling the model to learn contextualized representations from sequences of embedded tokens.
Patch Embedding. Patch embedding segments the input time series into patches every several timestamps along the temporal dimension using convolution, and projects each patch into a higher-dimensional space, similar to the patching mechanism in vision tasks dosovitskiy2020image . In the time-series context, patch embedding operates on each feature (channel) independently. The feed-forward model then treats the collection of patches as a batch, resulting in a channel-independent processing scheme, which is a design introduced alongside the patch embedding paradigm for time-series forecasting.
Invert Embedding. Invert embedding performs a linear projection across the temporal dimension for each feature independently, treating the entire lookback window of a single variable as a token. This representation is typically followed by a feed-forward model that focuses on learning inter-feature (token-wise) dependencies. It enables the model to capture relationships across different variables while maintaining the temporal integrity of each.
Frequency Embedding. Frequency embedding applies the RealFFT (rFFT) brigham2009fast to transform time-series sequences from the time domain into the frequency domain, enabling models to operate on spectral components rather than raw temporal signals. The resulting representation typically requires subsequent feed-forward models capable of handling complex-valued inputs, currently limited to MLPs in PyTorch implementations. To reconstruct the original temporal features, an inverse rFFT (irFFT) is employed in place of standard linear projections. Notably, frequency embedding is a non-parametric operation, distinguishing it from other learnable embedding methods.
2.3 Feed-Forward Modeling
In TimeRecipe, we consider various feed-forward modeling approaches from both a model architecture perspective (FF-type) and a model fusion perspective. Specifically, for model architectures, we include MLPs, Transformers, and RNNs. For model fusion strategies, we distinguish between temporal fusion, which aims to capture temporal dependencies, and feature fusion, which focuses on modeling correlations among features. We provide detailed descriptions of these feed-forward modeling approaches below.
Model Architectures. While Transformer-based models have achieved notable success in time-series forecasting, largely due to their ability to capture long-range dependencies across time steps or features, recent studies have also highlighted the effectiveness of MLP-based models for this task xu2023fits ; zeng2023transformers ; ni2025timedistill . Motivated by these findings, TimeRecipe aims to provide a deeper understanding of the relative effectiveness of different model architectures. For completeness, we also include RNNs as a baseline, representing a classical class of autoregressive models. RNN-based methods preceded the widespread adoption of Transformers, showing competitive performance in earlier work salinas2020deepar ; lai2018modeling and continue to demonstrate potential in more recent studies kong2025unlocking .
Modeling Fusion. Existing time-series forecasting approaches predict future values by either modeling temporal dependencies or capturing feature correlations. When considering fusion types across different model architectures, we highlight that the processing differs even under the same fusion type, depending on the underlying architecture, as shown below:
Temporal: $\text{MLP}(\text{X}_{t}^{\text{Emb}}\in[\text{B},\text{D},\text{L}]),\quad$ Feature: $\text{MLP}(\text{X}_{t}^{\text{Emb}}\in[\text{B},\text{L},\text{D}])$ (4)
$\text{RNN}(\text{X}_{t}^{\text{Emb}}\in[\text{B},\text{L},\text{D}]),$ $\text{RNN}(\text{X}_{t}^{\text{Emb}}\in[\text{B},\text{D},\text{L}])$
$\text{Transformer}(\text{X}_{t}^{\text{Emb}}\in[\text{B},\text{L},\text{D}]),$ $\text{Transformer}(\text{X}_{t}^{\text{Emb}}\in[\text{B},\text{D},\text{L}])$
Here, we slightly abuse the notations: B is the batchsize, L is the temporal dimension, and D is the feature dimension, which are the 1st, 2nd, and 3rd dimension after the embedding, respectively, as to generalize the different shapes produced by various embedding strategies.
2.4 TimeRecipe Benchmark
With the component designs introduced above, we now present TimeRecipe, a unified time-series forecasting framework that (i) implements the discussed module designs from each component of the canonical architecture, and (ii) automatically adjusts inter-module connections. TimeRecipe controls the use of different designs through hyperparameters, including whether to apply instance normalization and series decomposition, the choice of embedding type, model architecture, and fusion type. Accordingly, TimeRecipe constructs the full model pipeline by automatically adjusting hidden dimensions, initializing appropriate module connections, and applying proper tensor operations during the forward pass. Beyond providing a comprehensive benchmarking and intuition-driven study as in this work, TimeRecipe also offers practical benefits: it enables time-series researchers to conveniently build and evaluate models on their own data, and provides flexibility for exploring novel designs across different components within the canonical architecture framework.
Benchmark Scope and Coverage. Through exhaustive design space exploration enabled by TimeRecipe, the framework covers over 100 types of model architectures via module-level combinations. By adjusting the component configurations, TimeRecipe is able to encompass many popular time-series forecasting models, as partially illustrated in Table 1. Consequently, comprehensive benchmarking of TimeRecipe not only evaluates existing forecasting models, but also uncovers the potential of alternative combinations of model designs.
TimeRecipe Component Published Work
IN SD Fusion Embedding FF-Type Model Venue
Feature Invert Trans. iTransformer liu2023itransformer ICLR’24
Temporal Freq. MLP FITS xu2023fits ICLR’24
Temporal Patch Trans. PatchTST Yuqietal-2023-PatchTST ICLR’23
Temporal None MLP DLinear zeng2023transformers AAAI’23
Temporal Token Trans. Informer zhou2021informer AAAI’21
Table 1: Examples of existing methods that is covered by TimeRecipe benchmark.
Remark 1.
Channel-independence Yuqietal-2023-PatchTST ; han2024capacity is an important property in time-series forecasting models, referring to the ability to forecast multiple time series independently while sharing model parameters. TimeRecipe implicitly incorporates channel-independence in its evaluations by combining specific embedding types and model architectures with temporal modeling fusion. For example, using invert embedding with an MLP model architecture alongside temporal fusion serves as an instance of channel-independent modeling.
Remark 2.
We acknowledge that TimeRecipe does not cover every existing module design in time-series forecasting approaches. The uncovered designs generally fall into two categories. First, highly specialized designs, such as tangled temporal and feature fusion with delicate design used specifically in Crossformer zhang2023crossformer . Second, designs that are of less concern in terms of module effectiveness, such as data augmentation techniques based on down-sampling wang2024timemixer ; wang2024timemixer++ , which improve the forecasting accuracy over broad experiments.
3 Evaluation Protocol
3.1 Evaluation Tasks
We distinguish different time-series forecasting tasks from two perspectives. First, from a feature perspective, we categorize tasks into multivariate forecasting, where $d_{\text{X}}>1$ and the goal is to predict the future values of all feature dimensions simultaneously, and univariate forecasting, where $d_{\text{X}}=1$ and the task focuses on forecasting future values of a single dimension based solely on its own historical observations without exogenous features. Second, from a horizon length perspective, we classify tasks into short-term and long-term forecasting. Specifically, we define short-term forecasting as cases where the lookback window is longer than the forecasting horizon (i.e., $L>H$), and long-term forecasting as cases where the lookback window is shorter than or equal to the forecasting horizon (i.e., $L\leq H$).
Following this distinction, TimeRecipe aims to comprehensively benchmark the effectiveness of the modules described in Section 2 across all types of time-series forecasting tasks, including short-term univariate, short-term multivariate, long-term univariate, and long-term multivariate forecasting problems. These tasks broadly cover a wide range of time-series forecasting scenarios, where understanding the performance of different modules under varying task settings can offer practical insights for real-world forecasting applications.
3.2 Evaluation Dataset
LTSF. The Long-Term Time-Series Forecasting (LTSF) datasets zhou2021informer are widely adopted benchmarks for evaluating both short-term and long-term forecasting tasks. They test a model’s ability to generalize across diverse domains, including Electricity Transformer Temperature (ETT; ETTh1/2, ETTm1/2) zhou2021informer , Influenza-like Illness statistics (ILI), Electricity Consumption Load (ECLasuncion2007uci , meteorological data from the National Renewable Energy Laboratory (Weather), and foreign exchange rates across various countries (Exchange).
PEMS. The Performance Measurement System (PEMS, extended Traffic dataset in LTSF) datasets li2017diffusion are standard benchmarks for time-series forecasting, commonly used in traffic prediction research. These datasets contain road occupancy or flow measurements collected by loop detectors on highways across different districts in California. We include PEMS03, PEMS04, PEMS07, and PEMS08, each varying in geographic scope, number of sensors, and data volume.
M4. The M4 dataset makridakis2018m4 is a large-scale benchmark for evaluating forecasting models across diverse real-world time series. It includes 100,000 series from domains such as macroeconomics, microeconomics, finance, industry, and demography. Each time series varies in length and frequency, spanning yearly, quarterly, monthly, weekly, daily, and hourly settings.
We include additional details and the rationale behind the selection of these datasets in Appendix B.1.
3.3 Evaluation Metric
We adopt common evaluation metrics on each dataset. Specifically, we use mean squared error (MSE) and mean absolute error (MAE) for evaluations on LTSF and PEMS datasets, and symmetric mean absolute percentage error (SMAPE), mean absolute scaled error (MASE), and overall weighted average (OWA) for M4 dataset. The formula of these metrics is detailed in Appendix B.2.
Since error scales vary significantly across datasets, we use averaged rank scores to enable fair comparisons of module effectiveness. For example, if a module combination ranks 1st in MSE and 2nd in MAE, its average rank score is calculated as $1.5$. This unified ranking metric facilitates consistent evaluation across datasets and forms the basis of our analysis in Section 4.
3.4 Implementation and Reproducibility
Implementation. The implementation of TimeRecipe follows the widely-used Time-Series Library  [^1] https://github.com/thuml/Time-Series-Library wang2024deep . For usability, we introduce hyperparameters that control the inclusion of each module, allowing users to specify their usage of each module while TimeRecipe automatically adjusts the internal architecture and dimensions accordingly.
Reproducibility. For reproducibility, we provide all data, code, and scripts, publicly available at https://github.com/AdityaLab/TimeRecipe. All experiments are conducted on 32GB V100 GPUs. For statistical robustness, all results are based on the average over four random seeds.
4 Results and Discussion
In this section, we present the benchmark results along with key insights and discussions derived from them. Due to space constraints, we include partial results in Appendix C. and full results (both raw and processed) are available at: https://github.com/AdityaLab/TimeRecipeResults.
4.1 Insights Derived from Benchmark Results
4.1.1 TimeRecipe Yields Model Architectures Outperforming SOTA
A key observation from our evaluations is that exhaustive exploration of the design space across diverse datasets can identify model configurations that outperform existing SOTA approaches that are themselves included within the search space. For example, in the short-term multivariate forecasting on the PEMS03 dataset with a horizon of 12, the top-ranked configuration achieves an MSE of 0.714, outperforming iTransformer, one of the best existing forecasting models also covered by TimeRecipe, which attains an MSE of 0.739 and ranks only 7th among all evaluated design combinations.
More importantly, this phenomenon is not a rare case: it holds in 92 out of 102 evaluated scenarios, detailed in Appendix C.2, suggesting that in over 90% of cases, TimeRecipe can identify a model architecture that surpasses the SOTA. On average, the best existing approaches lag behind the top-performing configurations identified by TimeRecipe by 13.66 ranking positions. By exhaustively exploring the design space, TimeRecipe achieves an average forecasting error reduction of 5.4% compared to the best existing methods.
This observation underscores the importance of exhaustive design space search, as by TimeRecipe, for achieving superior forecasting performance across different scenarios. Moreover, TimeRecipe offers practical utility in the convenience of the construction of end-to-end forecasting pipelines. For instance, when introducing a novel embedding technique for time-series forecasting, researchers can leverage the canonical architecture of TimeRecipe to identify the optimal configuration that best complements the proposed component. This facilitates fair and effective evaluations across a wide range of scenarios and supports the development of new state-of-the-art solutions.
4.1.2 TimeRecipe Uncovers Modules Effectiveness Linking to Data Properties
The major motivation of TimeRecipe is to systematically investigate the effectiveness of various architectural modules across diverse time-series forecasting scenarios, especially in relation to the intrinsic characteristics of different datasets. This motivation stems from the variation observed in optimal architectural configurations across datasets. For instance, the top-performing models on multivariate forecasting tasks using the ETT datasets often combine patch embeddings with MLP or RNN-based feed-forward networks. In contrast, on the Electricity dataset, models that with invert embeddings with Transformer architectures yield the best results. Another noteworthy finding is that while instance normalization generally improves performance across many LTSF benchmarks, its performance on the PEMS datasets is degraded. These trends highlight the potential for uncovering meaningful interactions between model architecture choices and dataset properties.
To address this gap and answer our central research question: Which modules and model designs are most effective under specific time-series forecasting scenarios? As a first step, we establish a taxonomy of time-series data characteristics. This taxonomy includes key properties such as seasonality, trend, stationarity, transition, shifting, and correlation qiu2024tfb , as detailed in Appendix B.3. In addition to these intrinsic data properties, we further categorize forecasting tasks based on input structure—namely, whether the data is univariate or multivariate—as well as the number of input features (N-Feature) and the horizon-to-lookback ratio (HL-Ratio), which may also influence the effectiveness of specific module configurations.
Building on these data characteristics, we conduct a correlation analysis to explore the relationship between dataset properties and the effectiveness of various module designs. Specifically, we employ statistical hypothesis testing via t-tests to assess whether particular module configurations yield significantly improved performance under specific data conditions. This analysis directly supports our primary research goal of identifying the most appropriate design choices for diverse forecasting scenarios. A complete list of all statistically significant correlations (i.e., $p$-value $\leq 0.05$) is provided in Table 3, Appendix C.1. Through the analysis, we observe patterns among the statistically significant module-level effectiveness as follows:
Pre-processing. Time-series pre-processing techniques, such as instance normalization and series decomposition, consistently enhance ranking performance under complex conditions. These include high correlation, transition, seasonality, and shifting, as well as high HL-Ratio (i.e., long-term forecasting) and low stationarity. This suggests that pre-processing stabilizes input representations, especially for irregular time series.
Fusion. Temporal fusion is broadly effective in different settings. It improves performance when the n-feature is low, and under high HL-Ratio, correlation, transition, shifting, seasonality, and trend, highlighting the importance of capturing contextual temporal patterns in diverse forecasting scenarios.
Embedding Strategies. Embedding choices are more dependent on data properties. Invert embedding is favored when the n-feature, correlation, or transition is high. Token embedding is more effective under low HL-Ratio, trend, or stationarity. Patch embedding performs well in low-trend regimes. In relatively simple settings (e.g., low-trend univariate series), no embedding can also be beneficial.
Feedforward Modules. Transformer-based feedforward layers are favorable when correlation, seasonality, or transition is high, reflecting their strength in capturing global dependencies. Conversely, RNN-based feedforward layers perform better when n-feature count, HL-Ratio, shifting, trend, or stationarity is low, suggesting their suitability for simpler or more sequential patterns.
These findings suggest that architectural decisions should be tailored to the underlying data characteristics, with certain configurations consistently outperforming others under specific conditions.
4.2 Practical Implications from the Insights
Given the insights discussed above, a natural question arises: how can these findings be leveraged to inform better model architecture design for a given time-series dataset? To address this, TimeRecipe integrates a training-free model selection mechanism based on a LightGBM ke2017lightgbm regression model. Specifically, we train a regression model that maps both dataset characteristics and model configurations to their associated rank scores benchmarked by TimeRecipe. For a new forecasting task, we first compute the relevant data characteristics and then estimate the rank scores for a range of candidate configurations. The configuration with the lowest predicted rank score will be selected.
To assess the effectiveness of this training-free selection strategy, we evaluate it on two forecasting scenarios: (i) an in-distribution case involving short-term multivariate forecasting on the ETTh1 dataset (which is used in the TimeRecipe, though this specific setting is not benchmarked), and (ii) an out-of-distribution case using a new univariate unemployment (social) forecasting dataset introduced in Time-MMD liu2024timemmd , which is not part of the original TimeRecipe benchmark.
The results, presented in Table 2, demonstrate that even with a simple tree-based approach, TimeRecipe is capable of selecting models that are closer to the globally optimal architecture than the existing best-performing model. These findings underscore the potential of deploying TimeRecipe in real-world model selection scenarios, where it is able to make model selections that can surpass the SOTA performance without the need for training.
ETTh1_24_M
IN SD Fusion Embed FF-Type Rank MSE MAE
TimeRecipe Best Feature Patch MLP 1.5 0.2963 - 0.3467 -
Existing Best (PatchTST) Temporal Patch Trans. 4.0 0.2988 -0.9% 0.3520 -1.5%
Top-3 Selection Feature Patch MLP 1.5 0.2963 0.0% 0.3467 0.0%
Temporal Patch MLP 20.5 0.3096 -4.3% 0.3586 -3.4%
Temporal Patch Trans. 4.0 0.2988 -0.9% 0.3520 -1.5%
Selection better than existing best: Yes (Selection 1&3)
Social_12_S
IN SD Fusion Embed FF-Type Rank MSE MAE
TimeRecipe Best Feature Patch MLP 1.0 0.0854 - 0.2072 -
Existing Best (PatchTST) Temporal Patch Trans. 25.5 0.0994 -16.4% 0.2256 -8.9%
Top-3 Selection Feature Patch RNN 14.0 0.0950 -11.2% 0.2202 -6.3%
Temporal Invert RNN 5.5 0.0897 -5.0% 0.2123 -2.4%
Temporal Token MLP 70.0 0.1310 -53.4% 0.2689 -29.7%
Selection better than existing best: Yes (Selection 1&2)
Table 2: Comparison of model configurations selected by our method against the best-performing existing model and the global best results found through exhaustive design space search. On both ETTh1_24_M and Social_12_S, at least one of our top-3 selections consistently outperforms the existing best and closely approaches the optimal configuration. This demonstrates the effectiveness of our training-free model selection method, even when using a simple tree-based approach.
4.3 Implications for Future Research
We identify key benefits of TimeRecipe that can support future research in time-series forecasting.
Convenient Framework for Time-Series Module Design and Evaluation. The TimeRecipe’s canonical architecture provides a practical and extensible framework for testing new component designs in time-series forecasting. A key benefit is the ease with which novel modules can be integrated and evaluated in a standardized pipeline. For example, researchers can introduce new embedding strategies beyond those currently benchmarked and seamlessly insert them into the TimeRecipe framework. The system can then be used to automatically explore optimal configurations of the remaining components, such as pre-processing or feed-forward types, and thereby facilitate efficient and systematic evaluation of new design choices within a comprehensive forecasting pipeline.
Implications for Designing Advanced Time-Series Forecasting Models. TimeRecipe indicates that no single existing architecture consistently outperforms others across all time-series forecasting scenarios, even with taking recent SOTA into consideration shi2024time ; wang2024timemixer++ ; zhong2025time ; kong2025unlocking . However, rather than interpreting this as a limitation that precludes the development of broadly effective models, we argue that this insight points to the potential of hybrid architectures. Future SOTA models may benefit from integrating multiple architectural components tailored to different data properties. For example, ni2025timedistill has shown that some temporal patterns are better captured by Transformer-based models, while others are more effectively modeled by MLPs. Extending this intuition, TimeRecipe aims to reveal that similar complementarities exist across a wide range of design dimensions, such as normalization strategies, embedding schemes (e.g., patch vs. token), or architectural combinations. These findings suggest that next-generation forecasting systems, including foundation models, which are often derived from supervised learning approaches, could possibly achieve broader robustness and improved generalization by dynamically leveraging hybrid designs tailored to specific forecasting scenarios and pattern types.
Advancing AutoML for Time-Series Forecasting. Another important contribution of TimeRecipe lies in its potential to advance AutoML research in time-series forecasting. While earlier literature has explored AutoML for time series, these studies have predominantly focused on model-level searches combined with hyperparameter and learning scheme optimization alsharef2022time ; alsharef2022review ; shchur2023autogluon ; westergaard2024time . With the canonical architecture introduced by TimeRecipe, future AutoML for time-series forecasting may enable finer-grained, module-level search across different components of the forecasting pipeline. This approach allows AutoML systems to explore a significantly broader and more flexible design space, in conjunction with traditional hyperparameter and learning scheme tuning. As a result, TimeRecipe potentially benefits more granular and effective AutoML in time-series applications.
5 Conclusion
In this work, we introduced TimeRecipe, a benchmarking framework that systematically evaluates the effectiveness of individual modules for time-series forecasting. By decomposing forecast architectures into modular components, TimeRecipe allows a fine-grained analysis of design choices across a wide range of datasets and task settings. Our study presents comprehensive benchmark evaluations through more than 10,000 experiments to date, as well as the novel evaluation scope of module-level effectiveness to our best knowledge. We further include the limitation discussion in Appendix D.
Our results highlight two core insights: first, that exhaustive design space search results in model designs that can outperform SOTA; and second, that module effectiveness is highly data-dependent, with no single design universally superior across all scenarios. These findings not only challenge the notion of one-size-fits-all solutions but also motivate future research into adaptive and hybrid model architectures. By releasing our benchmark and toolkit, we aim to support the community in both evaluating existing designs and exploring new ones under a unified, interpretable framework.
6 Acknowledgment
This paper was supported in part by the NSF (Expeditions CCF-1918770, CAREER IIS-2028586, Medium IIS-1955883, Medium IIS-2106961, Medium IIS-2403240, PIPP CCF-2200269), CDC MInD program, Meta faculty gift, and funds/computing resources from Georgia Tech and GTRI.
Appendix A Related Works
Supervised Time-Series Forecasting. Time-series forecasting, predicting future values based on historical observations, is a long-standing challenge across numerous domains. Early approaches have been dominated by statistical methods. Autoregressive Integrated Moving Average (ARIMA) models [8] and Exponential Smoothing (ETS) techniques [25] have become foundational, modeling temporal dependencies and trends through established statistical principles. These methods often assume linearity or specific data structures and can struggle with complex, non-linear patterns present in many real-world datasets.
As a transformative architecture from deep learning, the introduction of the Transformer architecture [65], with its powerful self-attention mechanism, has marked a pivotal moment for time-series forecasting. Early efforts have focused on enhancing the efficiency and effectiveness of Transformers for long sequential data, a common characteristic of time series. Models like Reformer [29] have introduced techniques to reduce computational complexity, while Informer [86], Autoformer [75], Pyraformer [40], and FEDformer [87] have explored specialized attention mechanisms (e.g., sparse attention, auto-correlation, pyramidal attention) and decomposition strategies (e.g., trend-seasonal decomposition, frequency domain processing) to better capture temporal dependencies in long sequences. Further refinements have addressed specific challenges inherent in time-series data [35]. For instance, Non-stationary Transformers [44] have incorporated mechanisms to handle distribution shifts over time, ETSformer [72] has integrated principles from classical exponential smoothing, and Crossformer [82] has focused on modeling dependencies across different variates in multivariate settings. PatchTST [54] has proposed segmenting time series into patches, treating them as tokens, which proves highly effective for long-term forecasting. More recently, iTransformer [42] has inverted the roles of embedding and attention layers, achieving strong results.
Despite the success of Transformer variants, recent research has spurred a debate regarding their necessity, demonstrating that simpler architectures, particularly those based on Multi-Layer Perceptrons (MLPs) or even linear layers, can achieve competitive or superior performance on many benchmarks. DLinear [79] has proposed a simple linear model with decomposition, challenging the complexity of contemporary Transformers. Subsequently, various MLP-based models have emerged, often emphasizing efficiency and specialized designs. LightTS [81] has utilized sampling-oriented MLP structures, TSMixer [13] has employed an all-MLP architecture with mixing across time steps and features. TimeMixer [69] and TimeMixer++ [68] have further reflected on the competent duties of Transformer components and have repurposed the Transformer architecture.
Beyond Transformers and MLPs, other architectural paradigms continue to be explored. Convolutional approaches, such as MICN [66] using multi-scale local and global context modeling, SCINet [39] employing sample convolution and interaction and TimesNet [74] modeling temporal 2D variations, offer alternative ways to capture temporal features. Recurrent architectures are also being revisited, exhibiting strong performance for forecasting [30]. Recent explorations also include leveraging pretrained language models for forecasting tasks [26, 63, 36]. This diverse landscape, i.e., various model architectures, highlights the necessity and urgency of this time-series forecasting recipe at the model level, which is developed in this paper.
Time-Series Foundation Models. Time-series foundation models, pre-trained on vast amounts of diverse time-series data, aim to zero-shot adapt to unseen time-series datasets. Existing methods largely follow the architectural design of foundation language models. For instance, Chronos [5] is based on the T5 [57] architecture, TimesFM [16] employs the decoder-only Transformer architecture, Lag-Llama [59] explicitly leverages the Llama architecture. Recent advancements [41, 62] address computational scaling while enhancing capacity by introducing mixture of experts (MoE) [11] techniques from foundation language models. We envision that the model-level time-series forecasting recipe proposed in this paper provides key insights for the architectural design of the next generation of foundation time-series models.
Time-Series Benchmarks. Robust and standardized benchmarks are essential for evaluating the performance of time-series forecasting models, enabling fair comparison and reproducible research. Historically, the M4 and M5 competitions [49, 50] have served as crucial benchmarks, providing large collections of diverse time series primarily in business, demographic, and economic domains, along with rigorous evaluation protocols [49, 50]. Recent efforts focus on providing unified benchmarking frameworks and toolkits. TFB [56], for instance, introduces a scalable suite covering multiple domains and methods, emphasizing reproducibility and systematic, aspect-based analysis. Furthermore, BasicTS [34] focuses on benchmarking multivariate time series forecasting, ProbTS [80] benchmarks both point and distributional forecasting, and GIFT-Eval [2] is designed for benchmarking foundation time-series models. Recently, ReC4TS [38] benchmarks how different reasoning strategies enhance zero-shot time-series forecasting. In addition, Time-MMD [37] and CiK [6] are designed for benchmarking multimodal time-series forecasting. Different from all existing time-series benchmarks, our work uniquely focuses on evaluating the effectiveness of individual module-level design choices for time-series forecasting. To the best of our knowledge, this is the first benchmark to systematically address this underexplored yet critical aspect of model design.
Auto-ML in Time-Series. Automated Machine Learning (AutoML) aims to automate the end-to-end process of applying machine learning, enhancing accessibility and efficiency. AutoML research in time series can be categorized as: (1) automated feature engineering, which generates relevant temporal features like lags and rolling statistics [12, 15]; (2) automated model selection, searching across diverse model families from statistical methods (e.g., ARIMA) to machine learning (e.g., boosted trees) and deep learning architectures (e.g., LSTMs, Transformers) [78, 1, 61]; (3) hyperparameter optimization, using techniques like Bayesian optimization or evolutionary algorithms to tune model configurations [73, 22]; (4) neural architecture search, which specifically automates the design of deep learning model structures suitable for capturing complex temporal patterns [58, 76].
Appendix B Detailed Evaluation Setup
B.1 Additional Dataset Explanation
In this section, we will show more details of the datasets and our selection.
ETT, alias Electricity Transformer Temperature [86], contains collected oil temperature and electricity load data per minute of two Chinese stations between 2016/07 to 2018/07. The original data is then aggregated every one hour (ETTh1/2) or every 15 minutes (ETTm1/2).
ILI, alias Influenza-like Illness statistics, contains weekly influenza surveillance collected and released by the CDC  [^2] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html starting from 1997-98 influenza season.
ECL, alias Electricity Consumption Load [7], contains the hourly electricity consumption history of 321 clients covering two years.
Weather contains hourly meteorological data from 1,600 locations across the U.S. between 2010 and 2013.
Exchange [31] contains daily exchange rates of U.S. dollars in eight foreign countries, including Australia, Britain, Canada, Switzerland, China, Japan, New Zealand, and Singapore, between 1990 and 2016.
PEMS, alias Performance Measurement System, contains traffic data per minute from 325 sensors in the Bay Area collected by California Transportation Agencies (CalTrans) between January 2017 and May 2017. We adopted data from 4 sparsely located sensors in our evaluations.
M4 contains 100,000 domain-specific series from domains such as macroeconomics, microeconomics, finance, industry, and demography. Each time series is aggregated in various granularities, from hourly to yearly.
Our collected data covers multiple domains, multiple granularity, and multiple time ranges, ensuring plenty of diversity like the corresponding series. By evaluating such datasets, we guarantee that the outputted recipe from our method maintains considerable coverage in real-world time-series tasks.
B.2 Evaluation Metric Formulation
We follow the common evaluation metrics across diverse forecasting datasets [55, 39, 69]. We evaluate the LTSF dataset and PEMS dataset by two metrics: mean squared error (MSE) and mean absolute error (MAE), aiming to evaluate the errors under the same scale. Given a predicted sequence $\hat{\mathbf{Y}}=\{\hat{x}_{t+1},\dots,\hat{x}_{t+H}\}$ and corresponding ground truth $\mathbf{Y}=\{x_{t+1},\dots,x_{t+H}\}$ where $\mathbf{Y},\hat{\mathbf{Y}}\in\mathbb{R}^{H\times d_{X}}$, these two metrics are calculated by:
$\mathrm{MSE}$ $=\sum_{i=1}^{H}\frac{(x_{t+i}-\hat{x}_{t+i})^{2}}{H}$ (5)
$\mathrm{MAE}$ $=\sum_{i=1}^{H}\frac{|x_{t+i}-\hat{x}_{t+i}|}{H}$ (6)
Meanwhile, we evaluate the M4 dataset by three metrics: Symmetric Mean Absolute Percentage Error (SMAPE), Mean Absolute Scaled Error (MASE), and Overall Weighted Average (OWA), to normalize the errors across vast scales inside the dataset:
$\mathrm{SMAPE}$ $=\frac{200}{H}\sum_{i=1}^{H}\frac{|x_{t+i}-\hat{x}_{t+i}|}{|x_{t+i}|+|\hat{x}_{t+i}|}$ (7)
$\mathrm{MASE}$ $=\frac{1}{H}\sum_{i=1}^{H}\frac{|x_{t+i}-\hat{x}_{t+i}}{\frac{1}{H-S}\sum_{j=s+1}^{H}|x_{t+j}-x_{t+j-s}|}$ (8)
$\mathrm{OWA}$ $=\frac{1}{2}[\frac{\mathrm{SMAPE}}{\mathrm{SMAPE_{Naive2}}}+\frac{\mathrm{MASE}}{\mathrm{MASE_{Naive2}}}]$ (9)
B.3 Dataset Properties Measurement
We define the data properties involved in TimeRecipe following TFB [56].
Trend refers to the long-term changes of time-series. As shown in Algorithm 1, we involve Seasonal and Trend decomposition using Loess (STL) [14] as a decomposition function to calculate the trending values.
Seasonality refers to the changes in time-series that repeat every certain period. Similar to the calculation of Trend, we involve STL to calculate the seasonality, which is shown in Algorithm 2.
Stationarity refers to the indicator of whether a time-series approximately satisfies all of the following: 1) the mean of any observation inside the series is constant, 2) the variance of any observation is finite, 3) the covariance between any two observations depends only on their distance. The calculation is shown in Algorithm 3, where ADF refers to the Augmented Dickey-Fuller (ADF) test.
Shifting refers to time-series changes upon a certain direction over time. Given a threshold $m$, we calculate the shifting indicators in Algorithm 4.
Transition refers to the covariance of the transition matrix across symbols from 3-value windows. We calculate the transition values as shown in Algorithm 5,
Correlation refers to the possibility that different channels of a multivariate sequence share a similar distribution. As shown in algorithm 6, we calculate the correlation referring to Catch22 [47] library, which is designed to extract 22 features from time series.
Input : Time-Series $\mathbf{X}\in\mathbb{R}^{T\times 1}$
Output : Trend strength $\beta$ of $\mathbf{X}$
1
2S, T, R $\leftarrow$ STL(X);
3 return $\beta\leftarrow\max(0,1-\frac{\mathrm{var}(R)}{\mathrm{var}(T+R)})$;
Algorithm 1 Calculating Trend Strength of Time-Series
Input : Time-Series $\mathbf{X}\in\mathbb{R}^{T\times 1}$
Output : Seasonality strength $\zeta$ of $\mathbf{X}$
1 S, T, R $\leftarrow$ STL(X);
2 return $\zeta\leftarrow\max(0,1-\frac{\mathrm{var}(R)}{\mathrm{var}(S+R)})$;
Algorithm 2 Calculating Seasonality Strength of Time-Series
Input : Time-Series $\mathbf{X}\in\mathbb{R}^{T\times 1}$
Output : Stationarity value $\gamma$ of $\mathbf{X}$
1 $s\leftarrow\mathrm{ADF}(\mathbf{X})$;
2 return $\gamma\leftarrow(s\leq 0.05)$;
Algorithm 3 Calculating Stationarity Value of Time-Series
Input : Time-Series $\mathbf{X}\in\mathbb{R}^{T\times 1}$, number of thresholds $m$
Output : Shifting value $\delta$ of $\mathbf{X}$
1 $Z\leftarrow\mathrm{Z-normalize}(\mathbf{X}))$;
2 $Z_{min}\leftarrow\min(Z),\ Z_{max}\leftarrow\max(Z)$;
3 $S\leftarrow\{s_{i}|s_{i}\leftarrow Z_{min}+(i-1)\frac{Z_{max}-Z_{min}}{m},i\in[1,m]\}$;
4 for $i=1$ to $m$ do
5       $K_{i}\leftarrow\{j\,|\,Z_{j}>s_{i},j\in[1,T]\}$;
6       $M_{i}\leftarrow\mathrm{median}(K_{i})$
7 end for
[…]
9 return $\delta\leftarrow\mathrm{median}(M^{\prime})$;
Algorithm 4 Calculating Shifting Values of Time-Series
Input : Time-Series $\mathbf{X}\in\mathbb{R}^{T\times 1}$
Output : Transition value $\Delta\in(0,\frac{1}{3})$ of $\mathbf{X}$
1 $\tau\leftarrow$ First zero crossing of $X$’s autocorrelation;
2 $Y\leftarrow$ Downsampling $X$ with stride $\tau$;
3 $I\leftarrow\mathrm{argsort}(Y)$;
4 $T^{\prime}\leftarrow\mathrm{length}(Y)$ for $j\in[0:T^{\prime}]$ do
5       $Z_{j}\leftarrow\lfloor\frac{3I_{j}}{T^{\prime}}\rfloor$;
6
7 end for
[…]
9       $M[Z_{j}-1][Z_{j+1}-1]\leftarrow M[Z_{j}-1][Z_{j+1}-1]+1$
10 end for
[…]
12 $C\leftarrow$ covariance matrix between the columns of $M^{\prime}$;
return $\Delta\leftarrow tr(C)$
Algorithm 5 Calculating Transition Values of Time-Series
Input : Multivariate Time-Series $\mathbf{X}\in\mathbb{R}^{T\times N}$
Output : Correlation value $\eta\in(0,1)$ of $\mathbf{X}$
1 $F=\langle F^{1},\dots,F^{N}\rangle\in\mathcal{R}^{22\times N}\leftarrow\mathrm{Catch22}(\mathbf{X})$;
2 $P=\{r_{Pearson}(F^{i},F^{j})\,|\,i\in[1,N],j\in[i+1,N],i,j\in N^{*}\}$;
return $\eta\leftarrow\mathrm{mean}(P)+\frac{1}{1+\mathrm{var}(P)}$
Algorithm 6 Calculating Correlation Values of Time-Series
Appendix C Detailed Results
C.1 Comprehensive Claims (Section 4.1.2 Cont.)
Here we present all the claims on the correlations between the effectiveness of the module and the data characteristics that are statistically significant (e.g., $\leq$0.05). The comprehensive results are shown in Table 3.
Setting Property Direction Choice $p$-value
Multivariate N-Feature $\searrow$ IN=True 0.000
$\searrow$ Fusion=Temporal 0.000
$\nearrow$ Embed=Invert 0.000
$\nearrow$ FF=RNN 0.000
HL-Ratio $\searrow$ Embed=Token 0.002
$\searrow$ FF=RNN 0.000
$\nearrow$ IN=True 0.000
$\nearrow$ Fusion=Temporal 0.014
Correlation $\searrow$ IN=True 0.000
$\searrow$ Fusion=Temporal 0.000
$\nearrow$ Embed=Invert 0.000
$\nearrow$ FF=Transformer 0.000
Transition $\searrow$ Embed=Invert 0.012
$\searrow$ FF=Transformer 0.000
$\nearrow$ IN=True 0.000
$\nearrow$ Fusion=Temporal 0.000
Shifting $\searrow$ SD=True 0.034
$\searrow$ Embed=Invert 0.000
$\searrow$ FF=RNN 0.000
$\nearrow$ IN=True 0.000
$\nearrow$ Fusion=Temporal 0.000
Seasonality $\searrow$ IN=True 0.000
$\searrow$ Fusion=Temporal 0.000
$\nearrow$ Embed=Token 0.000
$\nearrow$ FF=Transformer 0.000
Trend $\searrow$ Embed=Token 0.006
$\searrow$ FF=RNN 0.000
$\nearrow$ IN=True 0.000
$\nearrow$ Fusion=Temporal 0.001
$\nearrow$ Embed=Patch 0.000
Stationarity $\searrow$ Embed=Token 0.021
$\searrow$ FF=RNN 0.000
$\nearrow$ Fusion=Temporal 0.026
Univariate HL-Ratio $\nearrow$ Fusion=Temporal 0.022
Transition $\nearrow$ SD=True 0.000
Shifting $\nearrow$ SD=True 0.000
$\nearrow$ Fusion=Temporal 0.011
Seasonality $\searrow$ Fusion=Temporal 0.008
$\nearrow$ FF=Transformer 0.000
Trend $\nearrow$ SD=True 0.000
$\nearrow$ Fusion=Temporal 0.000
$\nearrow$ Embed=None 0.044
Stationarity $\nearrow$ SD=True 0.002
Table 3: Statistical correlation between data/task properties and the effectiveness of specific architectural module choices in improving ranking performance (lower is better). Each row indicates whether a particular module configuration (e.g., input normalization, fusion type, embedding, feedforward block) is associated with a significantly lower or higher rank under certain data characteristics. The direction arrows ($\searrow$ or $\nearrow$) denote whether the choice is favored when the property is low or high, respectively, and the $p$-value quantifies the strength of the statistical association. Results are grouped by multivariate and univariate forecasting settings.
C.2 Performance Ranking Details (Section 4.1.1 Cont.)
To present a clearer comparison of the performance across different module combinations for each dataset and forecasting horizon, we present the top-10 ranked results. The ranking is based on the average of the ranks obtained from two metrics: MSE and MAE. Specifically, if a module combination ranks 1st in MSE and 2nd in MAE, its final rank score is computed as the average, i.e., 1.5. The detailed rankings are summarized in the following tables.
[…]
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ETTh1_96_M Feature Patch MLP 0.3761 1 0.3920 3 2.0
Feature Patch MLP 0.3783 3 0.3916 2 2.5
Temporal Patch RNN 0.3799 5 0.3952 5 5.0
Feature Patch MLP 0.3794 4 0.4011 8 6.0
Temporal Patch RNN 0.3803 7 0.3971 6 6.5
Temporal Frequency MLP 0.3776 2 0.4019 11 6.5
Temporal Patch RNN 0.3828 13 0.3915 1 7.0
Temporal None MLP 0.3801 6 0.4019 10 8.0
Temporal Patch RNN 0.3837 14 0.3930 4 9.0
Feature None Trans. 0.3811 9 0.4018 9 9.0
ETTh1_192_M Feature Patch MLP 0.4328 6 0.4238 2 4.0
Temporal Patch RNN 0.4314 4 0.4265 5 4.5
Temporal Patch RNN 0.4314 3 0.4285 6 4.5
Temporal Patch RNN 0.4339 9 0.4216 1 5.0
Temporal Frequency MLP 0.4306 2 0.4314 9 5.5
Temporal None MLP 0.4292 1 0.4318 10 5.5
Feature Patch MLP 0.4325 5 0.4310 8 6.5
Feature Patch MLP 0.4347 12 0.4260 4 8.0
Feature None Trans. 0.4336 8 0.4322 12 10.0
Temporal Patch MLP 0.4334 7 0.4343 17 12.0
ETTh1_336_M Temporal Invert MLP 0.4626 1 0.4462 2 1.5
Temporal Patch RNN 0.4766 7 0.4433 1 4.0
Temporal Patch Trans. 0.4713 3 0.4510 6 4.5
Temporal None MLP 0.4703 2 0.4526 10 6.0
Feature Patch MLP 0.4799 12 0.4467 3 7.5
Temporal Invert MLP 0.4774 8 0.4518 7 7.5
Temporal Patch RNN 0.4746 5 0.4541 14 9.5
Temporal Patch RNN 0.4835 16 0.4482 4 10.0
Temporal Frequency MLP 0.4744 4 0.4554 17 10.5
Feature None Trans. 0.4790 9 0.4538 13 11.0
ETTh1_720_M Temporal Invert MLP 0.4716 1 0.4698 1 1.0
Temporal Patch Trans. 0.4741 2 0.4722 3 2.5
Temporal Invert MLP 0.4818 4 0.4730 4 4.0
Temporal Patch RNN 0.4855 7 0.4719 2 4.5
Feature None Trans. 0.4792 3 0.4792 7 5.0
Temporal None MLP 0.4842 5 0.4787 6 5.5
Temporal Patch RNN 0.4853 6 0.4730 5 5.5
Feature Invert Trans. 0.4979 8 0.4825 9 8.5
Temporal None MLP 0.4982 10 0.4819 8 9.0
Temporal Frequency MLP 0.5009 12 0.4840 11 11.5
Table 4: Top-10 configurations for the ETTh1 dataset multivariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ETTh2_96_M Feature Patch MLP 0.2888 2 0.3392 1 1.5
Temporal Patch RNN 0.2895 3 0.3397 2 2.5
Temporal Patch RNN 0.2877 1 0.3422 5 3.0
Feature Patch MLP 0.2904 4 0.3406 3 3.5
Temporal Patch RNN 0.2917 7 0.3410 4 5.5
Temporal Invert MLP 0.2913 5 0.3432 9 7.0
Temporal Frequency MLP 0.2920 8 0.3426 6 7.0
Temporal None MLP 0.2915 6 0.3430 8 7.0
Feature Patch RNN 0.2944 11 0.3427 7 9.0
Temporal Frequency MLP 0.2927 9 0.3440 11 10.0
ETTh2_192_M Temporal Patch RNN 0.3684 2 0.3891 1 1.5
Feature Patch MLP 0.3677 1 0.3923 4 2.5
Feature Patch MLP 0.3705 3 0.3904 2 2.5
Temporal Patch MLP 0.3722 6 0.3917 3 4.5
Temporal Patch RNN 0.3718 5 0.3934 5 5.0
Temporal None MLP 0.3714 4 0.3952 11 7.5
Temporal Frequency MLP 0.3734 10 0.3937 6 8.0
Temporal None MLP 0.3723 7 0.3947 10 8.5
Feature Patch RNN 0.3731 8 0.3953 12 10.0
Temporal Invert MLP 0.3745 13 0.3943 8 10.5
ETTh2_336_M Temporal Patch RNN 0.4142 3 0.4269 1 2.0
Feature Invert Trans. 0.4131 2 0.4292 6 4.0
Temporal Invert MLP 0.4159 5 0.4286 3 4.0
Temporal Patch RNN 0.4164 6 0.4289 5 5.5
Feature None Trans. 0.4173 8 0.4286 4 6.0
Temporal Patch MLP 0.4158 4 0.4293 9 6.5
Feature Patch RNN 0.4166 7 0.4294 10 8.5
Temporal Patch MLP 0.4196 11 0.4292 7 9.0
Temporal Frequency MLP 0.4187 9 0.4306 11 10.0
Feature Patch MLP 0.4206 13 0.4292 8 10.5
Temporal Invert RNN 0.4269 21 0.4286 2 11.5
ETTh2_720_M Feature Invert Trans. 0.4218 1 0.4426 2 1.5
Temporal Invert MLP 0.4262 4 0.4426 1 2.5
Feature None Trans. 0.4257 3 0.4443 3 3.0
Temporal Patch MLP 0.4244 2 0.4446 5 3.5
Temporal Patch RNN 0.4265 5 0.4445 4 4.5
Temporal None MLP 0.4312 7 0.4455 6 6.5
Temporal None MLP 0.4290 6 0.4456 7 6.5
Temporal Patch MLP 0.4314 8 0.4483 10 9.0
Temporal Invert MLP 0.4320 10 0.4458 8 9.0
Temporal Invert RNN 0.4326 11 0.4464 9 10.0
Table 5: Top-10 configurations for the ETTh2 dataset multivariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ETTm1_96_M Feature Patch MLP 0.3274 1 0.3647 3 2.0
Temporal Patch RNN 0.3281 3 0.3640 2 2.5
Feature Patch RNN 0.3280 2 0.3654 4 3.0
Temporal Patch RNN 0.3308 8 0.3628 1 4.5
Temporal Patch MLP 0.3290 4 0.3661 7 5.5
Temporal None MLP 0.3294 6 0.3655 5 5.5
Temporal Invert RNN 0.3292 5 0.3669 9 7.0
Feature None Trans 0.3311 10 0.3659 6 8.0
Feature Patch MLP 0.3308 9 0.3662 8 8.5
Temporal Patch MLP 0.3299 7 0.3682 11 9.0
ETTm1_192_M Temporal Patch RNN 0.3681 3 0.3830 1 2.0
Feature Patch MLP 0.3669 2 0.3851 3 2.5
Temporal Patch RNN 0.3706 6 0.3836 2 4.0
Temporal Patch MLP 0.3667 1 0.3881 8 4.5
Temporal Patch MLP 0.3688 4 0.3872 6 5.0
Feature Patch RNN 0.3694 5 0.3862 5 5.0
Feature Patch MLP 0.3710 7 0.3855 4 5.5
Temporal None MLP 0.3722 8 0.3890 9 8.5
Temporal None MLP 0.3729 12 0.3878 7 9.5
Temporal Freq MLP 0.3724 9 0.3893 10 9.5
ETTm1_336_M Temporal Patch RNN 0.4001 3 0.4065 1 2.0
Feature Patch MLP 0.4003 4 0.4082 4 4.0
Temporal Patch MLP 0.4000 2 0.4100 7 4.5
Temporal Patch RNN 0.4025 7 0.4069 2 4.5
Temporal Patch MLP 0.3979 1 0.4109 9 5.0
Feature Patch RNN 0.4003 5 0.4093 5 5.0
Feature Patch MLP 0.4032 9 0.4073 3 6.0
Temporal Invert MLP 0.4030 8 0.4097 6 7.0
Temporal Freq MLP 0.4022 6 0.4107 8 7.0
Temporal None MLP 0.4058 12 0.4110 10 11.0
ETTm1_720_M Temporal Patch MLP 0.4577 2 0.4424 2 2.0
Temporal Patch MLP 0.4558 1 0.4441 3 2.0
Temporal Patch RNN 0.4628 4 0.4393 1 2.5
Temporal Patch MLP 0.4579 3 0.4486 10 6.5
Temporal Invert MLP 0.4653 7 0.4467 6 6.5
Temporal Patch RNN 0.4682 11 0.4462 4 7.5
Feature Patch MLP 0.4679 10 0.4463 5 7.5
Feature Patch RNN 0.4650 6 0.4485 9 7.5
Temporal Freq MLP 0.4648 5 0.4487 11 8.0
Temporal None MLP 0.4678 9 0.4484 8 8.5
Table 6: Top-10 configurations for the ETTm1 dataset multivariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ETTm2_96_M Temporal Invert RNN 0.1766 4 0.2576 1 2.5
Feature Patch MLP 0.1763 3 0.2581 3 3.0
Feature Patch RNN 0.1761 2 0.2590 4 3.0
Feature Patch MLP 0.1768 5 0.2577 2 3.5
Feature Patch RNN 0.1761 1 0.2600 9 5.0
Temporal Patch MLP 0.1771 7 0.2598 7 7.0
Temporal Freq MLP 0.1772 8 0.2596 6 7.0
Feature Token Trans. 0.1770 6 0.2599 8 7.0
Temporal Freq MLP 0.1795 16 0.2595 5 10.5
Feature Token Trans. 0.1783 11 0.2607 13 12.0
ETTm2_192_M Feature Patch MLP 0.2403 1 0.2996 1 1.0
Feature Patch MLP 0.2405 3 0.2997 2 2.5
Temporal Patch MLP 0.2405 2 0.3008 4 3.0
Temporal Invert RNN 0.2422 6 0.3008 5 5.5
Temporal Freq MLP 0.2406 4 0.3017 7 5.5
Temporal Patch RNN 0.2427 8 0.3008 6 7.0
Feature Patch RNN 0.2417 5 0.3024 9 7.0
Temporal Patch RNN 0.2432 11 0.3007 3 7.0
Temporal None MLP 0.2426 7 0.3023 8 7.5
Temporal Invert MLP 0.2428 9 0.3031 10 9.5
ETTm2_336_M Feature Patch MLP 0.2980 1 0.3378 2 1.5
Temporal Patch RNN 0.3009 5 0.3369 1 3.0
Feature Patch MLP 0.3002 3 0.3389 3 3.0
Temporal Freq MLP 0.2982 2 0.3393 5 3.5
Temporal Patch RNN 0.3017 8 0.3390 4 6.0
Temporal None MLP 0.3003 4 0.3405 8 6.0
Feature Patch RNN 0.3010 6 0.3398 6 6.0
Temporal Patch MLP 0.3013 7 0.3398 7 7.0
Temporal Freq MLP 0.3018 9 0.3405 9 9.0
Temporal Patch MLP 0.3024 10 0.3409 10 10.0
ETTm2_720_M Temporal Patch RNN 0.3942 1 0.3924 1 1.0
Feature Patch MLP 0.3981 3 0.3960 4 3.5
Temporal Invert MLP 0.3965 2 0.3961 5 3.5
Feature Patch MLP 0.3983 4 0.3958 3 3.5
Temporal Patch RNN 0.3991 8 0.3951 2 5.0
Temporal Patch MLP 0.3984 5 0.3963 6 5.5
Temporal Invert MLP 0.3990 6 0.3966 7 6.5
Temporal None MLP 0.3991 7 0.3968 8 7.5
Temporal Freq MLP 0.3992 9 0.3980 10 9.5
Temporal None MLP 0.4006 10 0.3975 9 9.5
Table 7: Top-10 configurations for the ETTm2 dataset multivariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
Exchange_96_M Feature Patch MLP 0.0821 1 0.1974 1 1.0
Temporal Patch RNN 0.0823 3 0.1978 2 2.5
Temporal Patch MLP 0.0823 2 0.1988 4 3.0
Temporal Patch RNN 0.0828 5 0.1987 3 4.0
Temporal Invert MLP 0.0829 6 0.1994 6 6.0
Feature Patch MLP 0.0833 7 0.1988 5 6.0
Temporal Patch Trans 0.0827 4 0.2002 8 6.0
Temporal Patch MLP 0.0836 9 0.2000 7 8.0
Temporal Invert MLP 0.0839 10 0.2004 9 9.5
Temporal None MLP 0.0845 11 0.2017 10 10.5
Exchange_192_M Temporal Patch MLP 0.1724 1 0.2939 1 1.0
Temporal None MLP 0.1734 2 0.2954 4 3.0
Feature Patch MLP 0.1742 4 0.2945 2 3.0
Temporal Patch RNN 0.1744 5 0.2948 3 4.0
Temporal Invert MLP 0.1741 3 0.2954 5 4.0
Temporal Patch RNN 0.1754 7 0.2958 6 6.5
Temporal Invert MLP 0.1751 6 0.2965 8 7.0
Feature Patch MLP 0.1757 9 0.2962 7 8.0
Feature Patch RNN 0.1757 8 0.2972 9 8.5
Temporal None MLP 0.1761 11 0.2979 11 11.0
Exchange_336_M Temporal Patch MLP 0.3192 1 0.4075 1 1.0
Temporal Invert MLP 0.3246 2 0.4121 2 2.0
Temporal Patch RNN 0.3278 5 0.4127 3 4.0
Feature Patch RNN 0.3278 3 0.4143 7 5.0
Feature Patch MLP 0.3279 6 0.4134 4 5.0
Temporal Patch RNN 0.3284 7 0.4135 5 6.0
Temporal None MLP 0.3278 4 0.4148 8 6.0
Feature Patch MLP 0.3296 8 0.4139 6 7.0
Feature Invert Trans. 0.3312 10 0.4165 9 9.5
Feature None Trans. 0.3307 9 0.4172 10 9.5
Exchange_720_M Temporal Freq MLP 0.8183 1 0.6837 1 1.0
Temporal None MLP 0.8337 3 0.6871 2 2.5
Temporal Patch MLP 0.8378 4 0.6872 3 3.5
Temporal Freq MLP 0.8205 2 0.6950 7 4.5
Temporal Patch Trans. 0.8411 5 0.6895 4 4.5
Temporal None MLP 0.8490 8 0.6955 9 8.5
Temporal Patch RNN 0.8544 11 0.6948 6 8.5
Feature Patch MLP 0.8545 12 0.6947 5 8.5
Feature Patch MLP 0.8542 10 0.6954 8 9.0
Temporal Invert MLP 0.8493 9 0.6958 10 9.5
Table 8: Top-10 configurations for the Exchange Rate dataset multivariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ILI_24_M Feature Patch Trans. 2.0338 2 0.8776 1 1.5
Feature Patch Trans. 2.0034 1 0.9064 3 2.0
Temporal Freq MLP 2.1586 3 0.8874 2 2.5
Temporal Invert MLP 2.1982 4 0.9084 4 4.0
Temporal None MLP 2.2416 5 0.9231 5 5.0
Temporal Token MLP 2.3280 9 0.9492 7 8.0
Temporal Token MLP 2.3096 7 0.9640 11 9.0
Temporal Freq MLP 2.4091 13 0.9563 9 11.0
Feature Invert RNN 2.4471 17 0.9480 6 11.5
Temporal Token Trans. 2.3071 6 1.0035 20 13.0
ILI_36_M Temporal Token MLP 1.9663 2 0.9067 1 1.5
Feature Invert RNN 1.9444 1 0.9104 2 1.5
Feature Invert Trans. 2.1458 3 0.9381 7 5.0
Temporal Invert MLP 2.1907 6 0.9234 5 5.5
Feature Invert Trans. 2.1581 4 0.9414 9 6.5
Feature Patch Trans. 2.2107 8 0.9257 6 7.0
Temporal Freq MLP 2.2220 11 0.9190 4 7.5
Feature Patch Trans. 2.2456 13 0.9117 3 8.0
Temporal Token MLP 2.1651 5 0.9516 12 8.5
Temporal Invert MLP 2.2197 10 0.9459 10 10.0
ILI_48_M Feature Patch Trans. 1.9918 2 0.8685 1 1.5
Temporal Token MLP 1.8707 1 0.8860 3 2.0
Temporal Freq MLP 2.0316 3 0.8844 2 2.5
Feature Invert Trans. 2.0593 5 0.8933 4 4.5
Temporal Freq MLP 2.0393 4 0.9152 8 6.0
Feature Invert Trans. 2.1405 7 0.9102 6 6.5
Feature Patch Trans. 2.1567 8 0.9074 5 6.5
Temporal Invert MLP 2.1383 6 0.9173 9 7.5
Feature Invert RNN 2.1660 9 0.9137 7 8.0
Temporal Token MLP 2.2079 12 0.9318 10 11.0
ILI_60_M Feature Patch Trans. 1.9860 1 0.8950 1 1.0
Feature Invert Trans. 2.0064 2 0.8998 2 2.0
Temporal Freq MLP 2.0119 3 0.9087 3 3.0
Feature Invert Trans. 2.0342 4 0.9130 4 4.0
Feature Invert RNN 2.0992 6 0.9336 5 5.5
Temporal Freq MLP 2.0898 5 0.9378 8 6.5
Temporal Invert MLP 2.1317 8 0.9346 6 7.0
Feature Patch Trans. 2.1360 9 0.9376 7 8.0
Feature Invert RNN 2.1096 7 0.9380 9 8.0
Temporal Patch MLP 2.1548 11 0.9435 10 10.5
Table 9: Top-10 configurations for the Illness (National Flu) dataset multivariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ECL_96_M Feature Invert Trans. 0.1545 1 0.2455 1 1.0
Feature Invert Trans. 0.1554 2 0.2523 3 2.5
Feature Invert Trans. 0.1599 3 0.2508 2 2.5
Feature Invert RNN 0.1602 4 0.2550 4 4.0
Feature Invert RNN 0.1618 5 0.2562 6 5.5
Feature None RNN 0.1630 7 0.2562 7 7.0
Feature Invert Trans. 0.1624 6 0.2578 9 7.5
Feature None Trans. 0.1682 12 0.2560 5 8.5
Feature None RNN 0.1651 10 0.2581 10 10.0
Feature Invert RNN 0.1634 8 0.2628 16 12.0
ECL_192_M Feature Invert Trans. 0.1670 1 0.2564 1 1.0
Feature Invert Trans. 0.1688 2 0.2640 3 2.5
Feature Invert Trans. 0.1707 3 0.2602 2 2.5
Feature Invert Trans. 0.1721 4 0.2663 5 4.5
Feature Invert RNN 0.1756 5 0.2682 10 7.5
Feature Patch Trans. 0.1802 12 0.2657 4 8.0
Feature Patch Trans. 0.1784 8 0.2676 9 8.5
Feature None RNN 0.1771 6 0.2683 11 8.5
Feature None Trans. 0.1797 11 0.2669 8 9.5
Feature Invert RNN 0.1772 7 0.2687 13 10.0
ECL_336_M Feature Invert Trans. 0.1827 1 0.2732 1 1.0
Feature Invert Trans. 0.1864 3 0.2769 2 2.5
Feature Invert Trans. 0.1852 2 0.2830 6 4.0
Feature Patch Trans. 0.1924 4 0.2835 7 5.5
Feature Patch Trans. 0.1965 11 0.2823 3 7.0
Feature Patch RNN 0.1942 7 0.2850 8 7.5
Feature Patch RNN 0.1947 8 0.2850 9 8.5
Feature Invert RNN 0.1937 6 0.2856 11 8.5
Feature Invert Trans. 0.1928 5 0.2873 15 10.0
Feature Patch RNN 0.1978 15 0.2825 5 10.0
ECL_720_M Feature Invert Trans. 0.2201 2 0.3061 1 1.5
Feature Invert Trans. 0.2221 3 0.3082 2 2.5
Feature Invert Trans. 0.2177 1 0.3153 5 3.0
Feature Patch Trans. 0.2275 5 0.3152 3 4.0
Feature Patch RNN 0.2299 6 0.3166 8 7.0
Feature Invert Trans. 0.2234 4 0.3176 11 7.5
Feature Patch RNN 0.2308 7 0.3167 9 8.0
Feature Patch Trans. 0.2364 14 0.3152 4 9.0
Feature Patch RNN 0.2385 17 0.3157 6 11.5
Feature None Trans. 0.2358 13 0.3168 10 11.5
Table 10: Top-10 configurations for the Electricity (ECL) dataset multivariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
Weather_96_M Feature Token RNN 0.1588 1 0.2062 1 1.0
Temporal Invert RNN 0.1589 2 0.2068 2 2.0
Feature Token RNN 0.1614 3 0.2092 3 3.0
Temporal Invert RNN 0.1617 4 0.2100 6 5.0
Feature Invert RNN 0.1634 7 0.2097 4 5.5
Feature None RNN 0.1634 6 0.2097 5 5.5
Temporal Token MLP 0.1624 5 0.2117 9 7.0
Feature None RNN 0.1642 9 0.2104 7 8.0
Feature Invert RNN 0.1649 13 0.2108 8 10.5
Temporal Token MLP 0.1648 11 0.2131 11 11.0
Weather_192_M Feature Token RNN 0.2096 3 0.2528 1 2.0
Feature Token RNN 0.2102 5 0.2531 2 3.5
Temporal Invert RNN 0.2100 4 0.2542 6 5.0
Feature None RNN 0.2119 10 0.2540 4 7.0
Feature Invert RNN 0.2126 13 0.2539 3 8.0
Feature Invert RNN 0.2127 14 0.2541 5 9.5
Feature None RNN 0.2125 12 0.2545 7 9.5
Temporal Invert RNN 0.2127 15 0.2558 9 12.0
Feature Patch RNN 0.2082 2 0.2656 38 20.0
Temporal Freq MLP 0.2058 1 0.2660 41 21.0
Weather_336_M Temporal Invert RNN 0.2671 14 0.2954 1 7.5
Feature Invert RNN 0.2700 18 0.2964 4 11.0
Feature Invert RNN 0.2704 20 0.2960 3 11.5
Feature None RNN 0.2705 22 0.2966 8 15.0
Temporal Freq MLP 0.2566 1 0.3054 34 17.5
Feature None RNN 0.2706 23 0.2973 14 18.5
Feature Token RNN 0.2719 26 0.2970 11 18.5
Temporal None MLP 0.2628 3 0.3078 39 21.0
Feature Patch RNN 0.2761 37 0.2965 7 22.0
Temporal Patch RNN 0.2631 5 0.3079 40 22.5
Weather_720_M Temporal Invert RNN 0.3429 19 0.3465 1 10.0
Feature Patch RNN 0.3536 38 0.3469 5 21.5
Temporal Patch MLP 0.3534 37 0.3471 7 22.0
Temporal Freq MLP 0.3548 41 0.3468 4 22.5
Temporal Token MLP 0.3481 24 0.3499 22 23.0
Temporal None MLP 0.3552 44 0.3466 2 23.0
Feature Invert RNN 0.3505 31 0.3485 15 23.0
Temporal Freq MLP 0.3549 42 0.3470 6 24.0
Feature Patch MLP 0.3355 10 0.3588 42 26.0
Feature Invert RNN 0.3517 32 0.3492 20 26.0
Table 11: Top-10 configurations for the Weather dataset multivariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
PEMS03_12_M Feature Invert RNN 0.0725 2 0.1777 1 1.5
Feature Invert Trans. 0.0714 1 0.1781 3 2.0
Feature Invert RNN 0.0726 3 0.1780 2 2.5
Feature None RNN 0.0726 4 0.1782 4 4.0
Feature None RNN 0.0734 5 0.1793 5 5.0
Feature Invert Trans. 0.0739 7 0.1812 6 6.5
Feature None Trans. 0.0739 6 0.1820 8 7.0
Feature Invert Trans. 0.0744 8 0.1814 7 7.5
Temporal Freq MLP 0.0750 10 0.1830 10 10.0
Temporal Invert RNN 0.0751 11 0.1823 9 10.0
PEMS03_24_M Feature Invert RNN 0.0985 1 0.2090 1 1.0
Temporal Invert RNN 0.0990 2 0.2091 2 2.0
Feature Invert RNN 0.1003 3 0.2108 3 3.0
Feature None RNN 0.1006 4 0.2125 5 4.5
Feature None RNN 0.1009 5 0.2117 4 4.5
Feature Token RNN 0.1023 6 0.2136 6 6.0
Temporal Invert RNN 0.1045 8 0.2140 7 7.5
Feature Invert Trans. 0.1030 7 0.2156 8 7.5
Feature Invert Trans. 0.1051 10 0.2156 9 9.5
Temporal Token MLP 0.1045 9 0.2176 11 10.0
PEMS03_36_M Feature Invert RNN 0.1220 1 0.2330 2 1.5
Temporal Invert RNN 0.1251 3 0.2315 1 2.0
Feature Invert RNN 0.1230 2 0.2348 4 3.0
Feature Token RNN 0.1254 4 0.2341 3 3.5
Temporal Invert RNN 0.1259 5 0.2371 5 5.0
Feature None RNN 0.1265 6 0.2386 6 6.0
Feature None RNN 0.1280 7 0.2393 7 7.0
Feature Token RNN 0.1294 8 0.2401 9 8.5
Feature Token RNN 0.1314 10 0.2395 8 9.0
Feature Token Trans. 0.1310 9 0.2410 11 10.0
PEMS03_48_M Feature Token RNN 0.1398 1 0.2489 1 1.0
Temporal Invert RNN 0.1426 3 0.2499 3 3.0
Temporal Token Trans. 0.1438 4 0.2530 5 4.5
Temporal Token RNN 0.1412 2 0.2543 7 4.5
Temporal Invert RNN 0.1461 9 0.2498 2 5.5
Feature Invert RNN 0.1445 5 0.2539 6 5.5
Feature Token RNN 0.1455 8 0.2525 4 6.0
Feature Invert RNN 0.1449 6 0.2552 8 7.0
Temporal Token Trans. 0.1450 7 0.2580 9 8.0
Feature Token Trans. 0.1499 11 0.2591 10 10.5
Table 12: Top-10 configurations for the PEMS03 dataset. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
PEMS04_12_M Feature Invert Trans. 0.0814 1 0.1891 1 1.0
Feature Invert RNN 0.0815 2 0.1894 2 2.0
Feature Invert RNN 0.0816 3 0.1899 3 3.0
Feature Invert Trans. 0.0825 4 0.1910 4 4.0
Feature None RNN 0.0826 5 0.1911 5 5.0
Feature None RNN 0.0829 6 0.1913 7 6.5
Temporal Invert RNN 0.0830 7 0.1912 6 6.5
Feature Token RNN 0.0856 9 0.1953 8 8.5
Temporal Invert RNN 0.0856 10 0.1955 9 9.5
Temporal Token MLP 0.0848 8 0.1958 11 9.5
PEMS04_24_M Temporal Invert RNN 0.0992 1 0.2119 1 1.0
Temporal Invert RNN 0.1013 2 0.2146 2 2.0
Temporal Token RNN 0.1013 3 0.2146 3 3.0
Temporal Token Trans. 0.1026 4 0.2151 6 5.0
Feature Invert RNN 0.1029 5 0.2168 10 7.5
Temporal Token Trans. 0.1043 8 0.2163 8 8.0
Feature Invert RNN 0.1032 6 0.2172 11 8.5
Temporal Invert RNN 0.1057 12 0.2155 7 9.5
Feature Token RNN 0.1034 7 0.2185 13 10.0
Feature Invert Trans. 0.1048 10 0.2187 14 12.0
PEMS04_36_M Temporal Invert RNN 0.1108 1 0.2203 1 1.0
Temporal Token RNN 0.1134 2 0.2208 2 2.0
Temporal Token Trans. 0.1149 4 0.2228 3 3.5
Temporal Token Trans. 0.1147 3 0.2230 4 3.5
Temporal Invert RNN 0.1155 6 0.2240 5 5.5
Temporal Invert RNN 0.1151 5 0.2294 8 6.5
Feature Token RNN 0.1157 7 0.2268 7 7.0
Temporal Token RNN 0.1184 10 0.2268 6 8.0
Temporal Token Trans. 0.1171 9 0.2318 12 10.5
Temporal Token RNN 0.1165 8 0.2323 14 11.0
PEMS04_48_M Temporal Invert RNN 0.1149 1 0.2246 1 1.0
Temporal Token Trans. 0.1168 3 0.2258 2 2.5
Temporal Token Trans. 0.1164 2 0.2261 3 2.5
Temporal Invert RNN 0.1174 4 0.2267 4 4.0
Temporal Token RNN 0.1195 6 0.2284 5 5.5
Temporal Token RNN 0.1192 5 0.2286 6 5.5
Feature Token RNN 0.1202 7 0.2312 7 7.0
Feature Invert MLP 0.1213 8 0.2322 8 8.0
Temporal None RNN 0.1236 9 0.2355 10 9.5
Temporal None RNN 0.1242 11 0.2334 9 10.0
Table 13: Top-10 configurations for the PEMS04 dataset. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
PEMS07_12_M Feature Invert Trans. 0.0663 1 0.1650 1 1.0
Feature Invert RNN 0.0664 2 0.1682 3 2.5
Feature Invert Trans. 0.0669 3 0.1676 2 2.5
Feature Invert RNN 0.0673 4 0.1696 5 4.5
Feature Invert Trans. 0.0682 7 0.1691 4 5.5
Feature None RNN 0.0676 5 0.1698 6 5.5
Feature None RNN 0.0677 6 0.1699 7 6.5
Feature Invert Trans. 0.0699 9 0.1712 8 8.5
Feature None Trans. 0.0697 8 0.1717 9 8.5
Feature Invert RNN 0.0703 10 0.1737 10 10.0
PEMS07_24_M Feature Invert Trans. 0.0901 1 0.1946 1 1.0
Feature Invert Trans. 0.0912 2 0.1959 2 2.0
Feature Invert RNN 0.0938 3 0.2009 3 3.0
Feature Invert RNN 0.0954 5 0.2027 5 5.0
Feature None RNN 0.0953 4 0.2031 6 5.0
Temporal Invert RNN 0.0999 8 0.2013 4 6.0
Feature None RNN 0.0962 6 0.2043 8 7.0
Temporal Invert RNN 0.1017 9 0.2037 7 8.0
Feature Invert Trans. 0.0992 7 0.2053 9 8.0
Feature Token RNN 0.1022 10 0.2057 10 10.0
PEMS07_36_M Feature Invert Trans. 0.1086 1 0.2159 1 1.0
Feature Invert Trans. 0.1118 2 0.2197 3 2.5
Temporal Invert RNN 0.1182 4 0.2195 2 3.0
Temporal Invert RNN 0.1206 6 0.2218 4 5.0
Feature Invert RNN 0.1175 3 0.2258 8 5.5
Feature Token RNN 0.1206 7 0.2234 5 6.0
Temporal Token RNN 0.1208 8 0.2241 6 7.0
Feature Invert RNN 0.1200 5 0.2282 12 8.5
Temporal Token RNN 0.1218 11 0.2249 7 9.0
Feature Invert MLP 0.1216 10 0.2259 9 9.5
PEMS07_48_M Feature Invert Trans. 0.1237 1 0.2319 1 1.0
Feature Invert Trans. 0.1283 2 0.2367 2 2.0
Temporal Invert RNN 0.1358 3 0.2378 3 3.0
Temporal Token RNN 0.1368 4 0.2396 4 4.0
Temporal Invert RNN 0.1377 8 0.2397 5 6.5
Temporal Token RNN 0.1373 6 0.2416 7 6.5
Feature Token RNN 0.1377 7 0.2399 6 6.5
Feature Invert RNN 0.1368 5 0.2457 10 7.5
Feature Invert RNN 0.1393 9 0.2472 14 11.5
Feature Invert MLP 0.1404 10 0.2472 13 11.5
Table 14: Top-10 configurations for the PEMS07 dataset. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
PEMS08_12_M Feature Invert Trans. 0.0795 1 0.1819 1 1.0
Feature Invert Trans. 0.0834 3 0.1868 2 2.5
Feature None Trans. 0.0833 2 0.1877 3 2.5
Feature Invert RNN 0.0844 4 0.1895 6 5.0
Feature Invert RNN 0.0856 5 0.1908 8 6.5
Feature None Trans. 0.0866 7 0.1913 10 8.5
Feature None RNN 0.0864 6 0.1918 11 8.5
Feature None RNN 0.0867 8 0.1920 12 10.0
Temporal Freq MLP 0.0870 9 0.1929 13 11.0
Temporal Freq MLP 0.0891 10 0.1955 14 12.0
PEMS08_24_M Feature Invert Trans. 0.1162 1 0.2191 1 1.0
Feature Invert Trans. 0.1226 2 0.2263 6 4.0
Feature None Trans. 0.1267 3 0.2327 8 5.5
Temporal Invert RNN 0.1318 6 0.2307 7 6.5
Feature Invert RNN 0.1296 4 0.2359 9 6.5
Feature Invert RNN 0.1302 5 0.2361 10 7.5
Temporal Freq MLP 0.1319 7 0.2382 12 9.5
Feature None RNN 0.1325 9 0.2384 13 11.0
Feature None RNN 0.1325 8 0.2386 14 11.0
Feature Token RNN 0.1378 14 0.2376 11 12.5
PEMS08_36_M Feature Invert Trans. 0.1581 1 0.2555 5 3.0
Temporal Invert RNN 0.1685 2 0.2638 7 4.5
Feature Invert Trans. 0.1689 3 0.2679 9 6.0
Feature Token RNN 0.1722 5 0.2681 10 7.5
Feature None Trans. 0.1711 4 0.2720 12 8.0
Temporal Invert RNN 0.1736 6 0.2710 11 8.5
Feature Token RNN 0.1745 7 0.2733 14 10.5
Temporal Token Trans. 0.1801 12 0.2725 13 12.5
Temporal Freq MLP 0.1768 8 0.2778 19 13.5
Feature Token Trans. 0.1769 9 0.2774 18 13.5
PEMS08_48_M Feature Invert RNN 0.2135 5 0.2645 1 3.0
Feature Invert Trans. 0.2013 1 0.2892 7 4.0
Temporal Invert RNN 0.2052 2 0.2947 10 6.0
Feature Invert RNN 0.2170 12 0.2673 2 7.0
Temporal Token Trans. 0.2112 4 0.2975 12 8.0
Feature Token RNN 0.2101 3 0.2986 13 8.0
Feature None RNN 0.2191 16 0.2714 3 9.5
Feature None RNN 0.2267 22 0.2788 5 13.5
Feature Token RNN 0.2161 9 0.3063 21 15.0
Feature Invert Trans. 0.2170 11 0.3049 19 15.0
Table 15: Top-10 configurations for the PEMS08 dataset. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type SMAPE Rank MASE Rank OWA Rank Total
M4_Hourly_S Feature Invert Trans. 18.0080 7 2.6195 3 1.0367 2 4.00
Temporal Invert MLP 17.8405 3 2.9283 5 1.0968 4 4.00
Feature Patch Trans. 18.2368 12 2.4683 1 1.0115 1 4.67
Temporal Patch MLP 17.6707 1 3.0920 8 1.1260 7 5.33
Temporal Invert MLP 17.9402 4 3.0478 7 1.1240 6 5.67
Temporal Patch MLP 17.8320 2 3.1867 9 1.1500 8 6.33
Temporal Patch Trans. 18.7640 20 2.5265 2 1.0377 3 8.33
Feature Patch RNN 18.0900 8 3.2920 12 1.1793 10 10.00
Feature Invert Trans. 19.0040 22 2.8540 4 1.1128 5 10.33
Temporal Patch RNN 17.9512 5 3.3385 15 1.1850 12 10.67
M4_Daily_S Temporal None MLP 2.9520 1 3.1000 1 0.9575 1 1.00
Temporal Freq MLP 2.9578 2 3.1227 2 0.9620 2 2.00
Feature Patch RNN 3.0127 4 3.1990 3 0.9825 3 3.33
Feature Patch RNN 3.0122 3 3.2135 4 0.9848 4 3.67
Feature Token RNN 3.0137 5 3.2145 5 0.9850 5 5.00
Temporal None MLP 3.0150 6 3.2190 7 0.9858 6 6.33
Temporal Invert MLP 3.0265 11 3.2172 6 0.9875 7 8.00
Temporal Patch MLP 3.0213 7 3.2287 10 0.9885 8 8.33
Temporal Freq MLP 3.0215 8 3.2322 12 0.9890 9 9.67
Feature Invert RNN 3.0270 12 3.2283 9 0.9892 11 10.67
M4_Weekly_S Feature Invert Trans. 9.3830 1 2.7800 1 1.0125 1 1.00
Feature Patch Trans. 9.4917 2 2.8413 2 1.0295 2 2.00
Feature Patch Trans. 9.5275 3 2.8433 3 1.0317 3 3.00
Feature Patch MLP 9.5322 4 2.9158 5 1.0453 4 4.33
Temporal Patch MLP 9.6020 6 2.9263 6 1.0507 5 5.67
Feature Invert Trans. 9.5980 5 2.9268 7 1.0508 6 6.00
Temporal Invert MLP 9.9285 12 2.8515 4 1.0550 7 7.67
Feature Patch MLP 9.6543 8 2.9610 10 1.0600 8 8.67
Temporal Patch MLP 9.6170 7 2.9830 11 1.0617 9 9.00
Feature Patch RNN 9.7745 9 2.9348 9 1.0620 10 9.33
M4_Monthly_S Feature Invert Trans. 12.5925 4 0.9193 1 0.8682 1 2.00
Temporal Freq MLP 12.5025 1 0.9273 6 0.8693 2 3.00
Temporal Freq MLP 12.5617 3 0.9253 4 0.8705 3 3.33
Feature Invert Trans. 12.6427 6 0.9195 2 0.8708 5 4.33
Temporal None MLP 12.5102 2 0.9290 10 0.8708 4 5.33
Feature Patch RNN 12.6077 5 0.9258 5 0.8725 7 5.67
Feature Token RNN 12.6585 8 0.9223 3 0.8723 6 5.67
Feature Token RNN 12.6908 15 0.9277 7 0.8760 8 10.00
Temporal Token MLP 12.6453 7 0.9335 13 0.8770 11 10.33
Feature Patch Trans. 12.6900 14 0.9280 8 0.8762 9 10.33
M4_Quarterly_S Temporal None MLP 9.9145 1 1.1500 1 0.8693 1 1.00
Temporal Freq MLP 9.9375 2 1.1542 2 0.8722 2 2.00
Temporal Freq MLP 10.0032 3 1.1687 5 0.8802 3 3.67
Feature Invert Trans. 10.0330 6 1.1667 4 0.8810 4 4.67
Temporal Freq MLP 10.0080 4 1.1700 7 0.8810 5 5.33
Feature Invert Trans. 10.0510 9 1.1667 3 0.8820 6 6.00
Temporal Token MLP 10.0237 5 1.1720 9 0.8825 8 7.33
Temporal Invert MLP 10.0347 7 1.1712 8 0.8825 7 7.33
Feature Patch Trans. 10.0590 12 1.1692 6 0.8830 9 9.00
Temporal Invert MLP 10.0503 8 1.1727 12 0.8840 11 10.33
M4_Yearly_S Temporal Freq MLP 13.2940 1 2.9985 2 0.7840 1 1.33
Temporal Freq MLP 13.3235 2 2.9930 1 0.7843 2 1.67
Temporal Invert MLP 13.3570 3 3.0072 3 0.7870 3 3.00
Temporal Token MLP 13.3673 5 3.0122 5 0.7880 5 5.00
Feature Patch MLP 13.3887 8 3.0077 4 0.7880 4 5.33
Temporal Freq MLP 13.3937 9 3.0202 11 0.7898 6 8.67
Temporal Invert RNN 13.4353 14 3.0130 6 0.7900 7 9.00
Temporal None MLP 13.3950 10 3.0230 13 0.7903 8 10.33
Temporal Freq MLP 13.3723 6 3.0380 21 0.7913 9 12.00
Temporal None MLP 13.3663 4 3.0393 23 0.7913 10 12.33
Table 16: Top-10 configurations for the M4 dataset. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ETTh1_96_S Temporal Patch RNN 0.0550 1 0.1781 1 1.0
Temporal None MLP 0.0550 2 0.1784 5 3.5
Feature Patch MLP 0.0552 5 0.1782 2 3.5
Feature None RNN 0.0552 4 0.1783 3 3.5
Temporal Patch RNN 0.0552 3 0.1785 6 4.5
Feature Invert MLP 0.0553 6 0.1785 7 6.5
Temporal Invert RNN 0.0554 10 0.1783 4 7.0
Temporal None MLP 0.0554 9 0.1785 8 8.5
Feature Invert RNN 0.0556 14 0.1787 9 11.5
Temporal Invert MLP 0.0555 11 0.1792 14 12.5
ETTh1_192_S Temporal Freq MLP 0.0716 1 0.2028 3 2.0
Temporal None MLP 0.0717 2 0.2027 2 2.0
Feature None RNN 0.0718 4 0.2026 1 2.5
Feature Invert RNN 0.0718 3 0.2030 4 3.5
Temporal Token MLP 0.0722 8 0.2035 7 7.5
Temporal Invert MLP 0.0720 6 0.2036 9 7.5
Feature Token RNN 0.0719 5 0.2038 12 8.5
Feature Token RNN 0.0722 9 0.2035 8 8.5
Feature Patch MLP 0.0725 14 0.2033 5 9.5
Temporal Patch RNN 0.0721 7 0.2039 13 10.0
ETTh1_336_S Temporal Freq MLP 0.0830 1 0.2242 1 1.0
Temporal None RNN 0.0833 3 0.2261 4 3.5
Temporal Invert RNN 0.0840 4 0.2263 6 5.0
Feature None RNN 0.0844 6 0.2261 5 5.5
Temporal Patch RNN 0.0842 5 0.2265 9 7.0
Feature Token RNN 0.0845 7 0.2264 7 7.0
Temporal None MLP 0.0850 13 0.2256 2 7.5
Temporal Token MLP 0.0849 12 0.2258 3 7.5
Temporal Patch MLP 0.0830 2 0.2270 13 7.5
Temporal Invert RNN 0.0845 8 0.2266 10 9.0
ETTh1_720_S Temporal Token RNN 0.0834 1 0.2284 2 1.5
Temporal Patch RNN 0.0841 2 0.2280 1 1.5
Feature Token MLP 0.0849 3 0.2302 3 3.0
Temporal Token Trans. 0.0851 4 0.2306 4 4.0
Temporal Patch Trans. 0.0862 5 0.2314 5 5.0
Temporal Patch RNN 0.0871 8 0.2317 7 7.5
Feature Token MLP 0.0867 6 0.2329 9 7.5
Feature Patch MLP 0.0871 9 0.2315 6 7.5
Feature None MLP 0.0871 7 0.2331 10 8.5
Temporal Invert RNN 0.0873 10 0.2321 8 9.0
Table 17: Top-10 configurations for the ETTh1 dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ETTh2_96_S Temporal Patch RNN 0.1275 1 0.2732 1 1.0
Temporal Patch RNN 0.1275 2 0.2734 2 2.0
Feature Patch MLP 0.1284 3 0.2747 3 3.0
Feature Invert MLP 0.1285 4 0.2747 4 4.0
Temporal Invert RNN 0.1287 5 0.2750 5 5.0
Feature None MLP 0.1290 6 0.2759 8 7.0
Temporal None MLP 0.1292 8 0.2756 6 7.0
Temporal Invert RNN 0.1293 9 0.2758 7 8.0
Feature Invert MLP 0.1292 7 0.2760 9 8.0
Feature Patch MLP 0.1296 10 0.2762 10 10.0
ETTh2_192_S Temporal Freq MLP 0.1680 1 0.3195 1 1.0
Feature Patch MLP 0.1780 4 0.3293 2 3.0
Temporal Patch MLP 0.1785 5 0.3315 6 5.5
Temporal Invert RNN 0.1797 11 0.3297 3 7.0
Temporal None MLP 0.1794 9 0.3315 7 8.0
Feature None MLP 0.1804 15 0.3309 4 9.5
Feature Invert MLP 0.1796 10 0.3325 11 10.5
Temporal None MLP 0.1787 7 0.3329 15 11.0
Temporal Patch RNN 0.1802 14 0.3316 8 11.0
Feature Invert MLP 0.1800 13 0.3319 9 11.0
ETTh2_336_S Feature Patch MLP 0.1933 2 0.3464 1 1.5
Feature Patch MLP 0.1908 1 0.3472 2 1.5
Temporal Freq MLP 0.1947 3 0.3507 3 3.0
Temporal Freq MLP 0.1969 5 0.3514 4 4.5
Temporal Patch Trans. 0.2005 8 0.3531 5 6.5
Temporal Token MLP 0.1969 4 0.3575 9 6.5
Feature Invert MLP 0.1984 6 0.3534 7 6.5
Feature Invert RNN 0.1988 7 0.3532 6 6.5
Feature Invert Trans. 0.2025 9 0.3564 8 8.5
Feature Invert Trans. 0.2051 11 0.3586 10 10.5
ETTh2_720_S Temporal Patch Trans. 0.2214 1 0.3782 1 1.0
Feature Patch MLP 0.2217 2 0.3790 2 2.0
Temporal Invert MLP 0.2238 3 0.3806 3 3.0
Feature Invert Trans. 0.2257 5 0.3814 4 4.5
Feature Invert RNN 0.2256 4 0.3823 5 4.5
Feature Patch MLP 0.2260 6 0.3828 6 6.0
Temporal Token MLP 0.2267 7 0.3834 8 7.5
Temporal None MLP 0.2269 8 0.3833 7 7.5
Feature Invert RNN 0.2275 9 0.3840 9 9.0
Temporal Patch MLP 0.2297 11 0.3847 10 10.5
Table 18: Top-10 configurations for the ETTh2 dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ETTm1_96_S Temporal Patch RNN 0.0286 1 0.1255 1 1.0
Feature Patch MLP 0.0287 2 0.1258 3 2.5
Feature Invert MLP 0.0287 3 0.1256 2 2.5
Feature Patch RNN 0.0287 4 0.1260 4 4.0
Temporal Invert RNN 0.0288 6 0.1260 5 5.5
Feature None RNN 0.0288 5 0.1260 6 5.5
Feature Invert RNN 0.0288 8 0.1261 8 8.0
Temporal Invert RNN 0.0288 7 0.1263 11 9.0
Feature None MLP 0.0289 10 0.1263 12 11.0
Temporal Invert MLP 0.0289 13 0.1262 10 11.5
ETTm1_192_S Temporal None RNN 0.0433 1 0.1583 3 2.0
Feature None MLP 0.0436 3 0.1582 1 2.0
Feature None MLP 0.0435 2 0.1585 4 3.0
Feature Invert MLP 0.0437 5 0.1582 2 3.5
Feature Patch RNN 0.0436 4 0.1586 6 5.0
Temporal Patch MLP 0.0438 7 0.1586 5 6.0
Feature Token MLP 0.0437 6 0.1587 7 6.5
Temporal Invert RNN 0.0439 9 0.1588 8 8.5
Temporal Patch MLP 0.0439 10 0.1588 9 9.5
Temporal Patch RNN 0.0440 11 0.1588 10 10.5
ETTm1_336_S Feature None MLP 0.0568 1 0.1837 2 1.5
Feature None MLP 0.0572 2 0.1831 1 1.5
Temporal Patch MLP 0.0574 3 0.1840 4 3.5
Feature Patch RNN 0.0576 5 0.1839 3 4.0
Feature Token MLP 0.0575 4 0.1843 7 5.5
Temporal None MLP 0.0577 7 0.1840 5 6.0
Feature Invert MLP 0.0578 8 0.1841 6 7.0
Temporal Invert RNN 0.0579 9 0.1845 9 9.0
Temporal None RNN 0.0576 6 0.1848 12 9.0
Temporal Freq MLP 0.0581 11 0.1845 8 9.5
ETTm1_720_S Temporal Invert RNN 0.0809 2 0.2172 1 1.5
Feature None MLP 0.0807 1 0.2175 2 1.5
Feature Freq MLP 0.0810 3 0.2179 5 4.0
Temporal None MLP 0.0811 4 0.2179 6 5.0
Feature Token MLP 0.0811 5 0.2180 9 7.0
Feature Invert RNN 0.0815 11 0.2177 3 7.0
Temporal Freq MLP 0.0815 12 0.2177 4 8.0
Temporal Token MLP 0.0815 9 0.2182 10 9.5
Temporal Patch MLP 0.0816 13 0.2179 7 10.0
Temporal Patch MLP 0.0812 6 0.2185 17 11.5
Table 19: Top-10 configurations for the ETTm1 dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ETTm2_96_S Feature None RNN 0.0655 1 0.1823 1 1.0
Feature Patch MLP 0.0658 5 0.1828 2 3.5
Feature Invert RNN 0.0657 3 0.1834 4 3.5
Temporal Invert RNN 0.0658 6 0.1831 3 4.5
Feature Invert Trans 0.0657 4 0.1838 8 6.0
Feature None RNN 0.0660 7 0.1836 5 6.0
Temporal Patch MLP 0.0660 8 0.1836 7 7.5
Feature Patch MLP 0.0661 9 0.1836 6 7.5
Temporal Freq MLP 0.0657 2 0.1848 14 8.0
Temporal Invert RNN 0.0665 13 0.1844 10 11.5
ETTm2_192_S Feature Patch RNN 0.0989 1 0.2325 1 1.0
Temporal Invert RNN 0.0996 3 0.2332 2 2.5
Feature Patch MLP 0.0996 5 0.2335 3 4.0
Feature Invert MLP 0.0996 4 0.2338 5 4.5
Feature None RNN 0.0998 7 0.2337 4 5.5
Feature None RNN 0.0997 6 0.2341 7 6.5
Temporal None MLP 0.0995 2 0.2347 12 7.0
Temporal Patch MLP 0.1005 13 0.2340 6 9.5
Feature Invert RNN 0.1000 9 0.2345 10 9.5
Temporal Patch MLP 0.1005 14 0.2344 8 11.0
ETTm2_336_S Feature None MLP 0.1289 1 0.2723 1 1.0
Feature None MLP 0.1290 2 0.2728 2 2.0
Feature Invert MLP 0.1297 3 0.2732 3 3.0
Feature Patch MLP 0.1304 6 0.2736 4 5.0
Temporal Token MLP 0.1303 5 0.2744 7 6.0
Temporal Patch RNN 0.1308 8 0.2743 6 7.0
Temporal Token MLP 0.1302 4 0.2746 12 8.0
Temporal Patch RNN 0.1311 13 0.2740 5 9.0
Temporal Invert RNN 0.1308 7 0.2745 11 9.0
Temporal Invert RNN 0.1309 9 0.2744 9 9.0
ETTm2_720_S Feature None MLP 0.1795 1 0.3290 2 1.5
Feature None MLP 0.1802 2 0.3287 1 1.5
Temporal Patch MLP 0.1824 5 0.3306 3 4.0
Temporal None RNN 0.1822 3 0.3315 5 4.0
Feature Invert MLP 0.1823 4 0.3312 4 4.0
Temporal Token MLP 0.1827 6 0.3318 7 6.5
Temporal Invert RNN 0.1829 8 0.3318 6 7.0
Temporal Freq MLP 0.1830 9 0.3320 9 9.0
Temporal Freq MLP 0.1828 7 0.3322 13 10.0
Temporal Invert RNN 0.1833 11 0.3322 11 11.0
Table 20: Top-10 configurations for the ETTm2 dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
Exchange_96_S Temporal Patch RNN 0.0887 1 0.2202 1 1.0
Temporal Patch MLP 0.0923 2 0.2230 2 2.0
Temporal Token MLP 0.0925 3 0.2246 4 3.5
Feature Patch MLP 0.0942 4 0.2237 3 3.5
Feature Invert MLP 0.0974 7 0.2283 6 6.5
Temporal Freq MLP 0.0971 6 0.2285 8 7.0
Temporal Invert RNN 0.0969 5 0.2298 10 7.5
Feature Patch MLP 0.0995 11 0.2281 5 8.0
Temporal None MLP 0.0978 8 0.2293 9 8.5
Feature None RNN 0.0986 10 0.2299 11 10.5
Exchange_192_S Temporal Patch MLP 0.1998 1 0.3336 1 1.0
Temporal Patch RNN 0.2047 4 0.3353 2 3.0
Temporal Invert RNN 0.2046 3 0.3356 5 4.0
Temporal Invert RNN 0.2056 7 0.3355 4 5.5
Temporal Token MLP 0.2038 2 0.3390 11 6.5
Feature None RNN 0.2068 9 0.3369 6 7.5
Feature Patch MLP 0.2082 15 0.3354 3 9.0
Feature Invert MLP 0.2053 5 0.3393 13 9.0
Feature Patch MLP 0.2072 12 0.3370 7 9.5
Temporal Invert MLP 0.2055 6 0.3404 14 10.0
Exchange_336_S Feature None Trans 0.4017 1 0.4832 1 1.0
Feature None RNN 0.4168 4 0.4854 5 4.5
Feature Patch RNN 0.4204 8 0.4852 3 5.5
Feature Patch RNN 0.4189 6 0.4865 6 6.0
Temporal Patch RNN 0.4217 11 0.4841 2 6.5
Feature None RNN 0.4205 9 0.4853 4 6.5
Temporal None MLP 0.4166 2 0.4898 13 7.5
Temporal Token MLP 0.4182 5 0.4901 17 11.0
Temporal Patch MLP 0.4247 15 0.4872 8 11.5
Feature Invert RNN 0.4247 16 0.4880 10 13.0
Exchange_720_S Temporal Token RNN 0.7083 1 0.7244 1 1.0
Temporal Invert MLP 0.8477 3 0.7296 2 2.5
Feature Token MLP 0.8456 2 0.7729 5 3.5
Temporal Invert RNN 0.8950 5 0.7391 3 4.0
Temporal Freq MLP 0.9377 6 0.7404 4 5.0
Feature Token MLP 0.8575 4 0.7779 6 5.0
Feature None Trans 1.0907 11 0.7944 8 9.5
Temporal None MLP 1.0925 14 0.7931 7 10.5
Feature None MLP 1.0923 13 0.7972 9 11.0
Feature Token Trans 1.0908 12 0.8024 11 11.5
Table 21: Top-10 configurations for the Exchange Rate dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ILI_24_S Temporal Patch RNN 0.6131 1 0.5881 3 2.0
Feature Patch Trans. 0.6292 2 0.5877 2 2.0
Temporal Patch MLP 0.6405 4 0.5831 1 2.5
Temporal Invert RNN 0.6338 3 0.5981 6 4.5
Feature Patch RNN 0.6613 7 0.5920 4 5.5
Feature None Trans. 0.6568 6 0.5949 5 5.5
Feature None Trans. 0.6564 5 0.6093 12 8.5
Temporal Patch RNN 0.6904 10 0.6036 8 9.0
Temporal Token MLP 0.6968 12 0.6036 7 9.5
Feature Invert MLP 0.6791 9 0.6081 10 9.5
ILI_36_S Temporal Patch MLP 0.6331 1 0.6096 1 1.0
Feature Patch Trans 0.6416 2 0.6118 2 2.0
Temporal Token RNN 0.6491 3 0.6270 3 3.0
Feature Invert MLP 0.6713 6 0.6312 5 5.5
Temporal Patch RNN 0.6636 4 0.6345 8 6.0
Feature Patch RNN 0.6790 8 0.6302 4 6.0
Temporal Patch RNN 0.6649 5 0.6381 10 7.5
Feature Patch MLP 0.7038 12 0.6321 6 9.0
Temporal Invert RNN 0.6750 7 0.6462 12 9.5
Feature Patch Trans. 0.7069 14 0.6327 7 10.5
ILI_48_S Feature Patch Trans. 0.6473 1 0.6268 1 1.0
Temporal Patch MLP 0.6568 2 0.6412 2 2.0
Feature Invert Trans. 0.6850 4 0.6439 3 3.5
Feature Patch RNN 0.6836 3 0.6539 5 4.0
Feature Patch RNN 0.7004 6 0.6603 8 7.0
Feature Patch Trans. 0.7013 8 0.6584 7 7.5
Temporal Token RNN 0.6879 5 0.6635 10 7.5
Temporal Invert MLP 0.7047 12 0.6622 9 10.5
Temporal Invert MLP 0.7113 18 0.6645 11 14.5
Feature Token RNN 0.7095 15 0.6671 14 14.5
ILI_60_S Feature Patch Trans. 0.6730 1 0.6682 1 1.0
Feature Invert Trans. 0.6822 2 0.6739 3 2.5
Temporal Patch MLP 0.6992 4 0.6830 7 5.5
Feature Patch Trans. 0.7022 6 0.6784 5 5.5
Temporal Patch MLP 0.6969 3 0.6849 9 6.0
Temporal Patch Trans. 0.7154 13 0.6744 4 8.5
Temporal Invert MLP 0.7008 5 0.6883 13 9.0
Temporal Patch Trans. 0.7215 16 0.6715 2 9.0
Feature Invert Trans. 0.7085 8 0.6872 11 9.5
Temporal Invert MLP 0.7113 10 0.6866 10 10.0
Table 22: Top-10 configurations for the Illness (National Flu) dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
ECL_96_S Feature Invert Trans. 0.2829 2 0.3794 2 2.0
Feature None Trans. 0.2825 1 0.3798 3 2.0
Feature Invert Trans. 0.2899 5 0.3768 1 3.0
Feature Invert MLP 0.2879 4 0.3829 4 4.0
Feature Invert MLP 0.2868 3 0.3833 5 4.0
Feature Invert Trans. 0.2946 7 0.3848 7 7.0
Feature None Trans. 0.2930 6 0.3852 9 7.5
Feature None RNN 0.2973 11 0.3838 6 8.5
Feature Patch RNN 0.2962 10 0.3852 8 9.0
Feature Invert MLP 0.2947 8 0.3899 14 11.0
ECL_192_S Feature Invert MLP 0.3013 1 0.3917 2 1.5
Feature None Trans. 0.3019 2 0.3912 1 1.5
Feature Invert MLP 0.3107 3 0.3961 5 4.0
Feature Invert Trans. 0.3128 5 0.3931 3 4.0
Feature None Trans. 0.3112 4 0.3957 4 4.0
Feature Invert Trans. 0.3145 6 0.4006 9 7.5
Feature Invert Trans. 0.3199 9 0.3962 6 7.5
Feature Invert MLP 0.3193 8 0.4007 10 9.0
Temporal Patch RNN 0.3223 11 0.4009 11 11.0
Temporal Invert RNN 0.3183 7 0.4043 20 13.5
ECL_336_S Feature Invert MLP 0.3437 1 0.4245 3 2.0
Feature Invert Trans. 0.3452 2 0.4243 2 2.0
Feature None Trans. 0.3462 3 0.4250 4 3.5
Feature Invert Trans. 0.3575 7 0.4226 1 4.0
Temporal Invert RNN 0.3489 4 0.4277 6 5.0
Feature None Trans. 0.3554 6 0.4270 5 5.5
Feature Invert MLP 0.3546 5 0.4287 8 6.5
Feature Invert MLP 0.3665 9 0.4286 7 8.0
Feature Patch RNN 0.3578 8 0.4295 9 8.5
Feature Patch Trans. 0.3665 10 0.4327 13 11.5
ECL_720_S Feature Invert MLP 0.3867 1 0.4638 1 1.0
Feature Invert MLP 0.3997 3 0.4682 3 3.0
Feature None Trans. 0.3974 2 0.4691 4 3.0
Feature Invert Trans. 0.4095 6 0.4646 2 4.0
Feature Invert Trans. 0.4067 4 0.4745 7 5.5
Feature None Trans. 0.4167 10 0.4706 5 7.5
Temporal Invert RNN 0.4112 7 0.4756 9 8.0
Feature Freq MLP 0.4076 5 0.4764 11 8.0
Feature Patch MLP 0.4165 9 0.4756 8 8.5
Temporal Patch RNN 0.4237 12 0.4724 6 9.0
Table 23: Top-10 configurations for the Electricity (ECL) dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
Weather_96_S Temporal Patch RNN 0.0012 1 0.0254 1 1.0
Temporal Patch RNN 0.0012 2 0.0255 2 2.0
Temporal Invert RNN 0.0013 3 0.0258 4 3.5
Temporal Invert RNN 0.0013 4 0.0260 5 4.5
Feature Patch RNN 0.0013 6 0.0256 3 4.5
Feature None MLP 0.0013 5 0.0261 7 6.0
Feature Patch MLP 0.0013 8 0.0261 6 7.0
Feature None MLP 0.0013 7 0.0263 8 7.5
Feature Patch RNN 0.0013 9 0.0264 10 9.5
Feature Patch MLP 0.0013 11 0.0264 9 10.0
Weather_192_S Temporal Token RNN 0.0015 1 0.0280 1 1.0
Temporal Token Trans. 0.0015 2 0.0286 3 2.5
Temporal Token RNN 0.0015 3 0.0287 6 4.5
Temporal Freq MLP 0.0015 7 0.0286 2 4.5
Temporal Patch RNN 0.0015 5 0.0287 5 5.0
Temporal Invert RNN 0.0015 4 0.0288 7 5.5
Feature Patch RNN 0.0015 9 0.0286 4 6.5
Feature None MLP 0.0015 6 0.0288 8 7.0
Feature Token MLP 0.0015 8 0.0291 11 9.5
Feature Patch RNN 0.0015 10 0.0291 13 11.5
Weather_336_S Feature Patch RNN 0.0017 4 0.0300 1 2.5
Temporal Token RNN 0.0016 1 0.0302 6 3.5
Temporal Token Trans. 0.0017 6 0.0301 2 4.0
Temporal Patch RNN 0.0016 2 0.0303 9 5.5
Feature None MLP 0.0017 7 0.0301 4 5.5
Temporal None MLP 0.0017 3 0.0304 11 7.0
Temporal Freq MLP 0.0017 13 0.0301 3 8.0
Feature Token MLP 0.0017 11 0.0303 7 9.0
Feature Invert MLP 0.0017 5 0.0307 13 9.0
Temporal Freq MLP 0.0017 15 0.0302 5 10.0
Weather_720_S Temporal Token Trans. 0.0021 2 0.0335 1 1.5
Temporal Freq MLP 0.0021 1 0.0337 2 1.5
Temporal Freq MLP 0.0021 3 0.0337 3 3.0
Feature Patch RNN 0.0021 4 0.0341 5 4.5
Feature Token MLP 0.0021 6 0.0340 4 5.0
Feature None MLP 0.0021 10 0.0341 6 8.0
Temporal None MLP 0.0021 5 0.0344 11 8.0
Temporal Token RNN 0.0021 7 0.0346 13 10.0
Feature Token MLP 0.0022 14 0.0341 7 10.5
Temporal None MLP 0.0021 9 0.0347 14 11.5
Table 24: Top-10 configurations for the Weather dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
PEMS03_12_S Temporal Freq MLP 0.0325 1 0.1331 1 1.0
Temporal Freq MLP 0.0333 2 0.1340 2 2.0
Temporal Token MLP 0.0346 3 0.1357 4 3.5
Temporal Token MLP 0.0346 4 0.1353 3 3.5
Feature Token RNN 0.0350 5 0.1368 5 5.0
Temporal Token RNN 0.0350 6 0.1371 6 6.0
Feature Patch RNN 0.0354 7 0.1381 8 7.5
Temporal None MLP 0.0359 11 0.1378 7 9.0
Feature Token RNN 0.0358 9 0.1383 9 9.0
Feature Patch RNN 0.0358 10 0.1384 10 10.0
PEMS03_24_S Temporal Freq MLP 0.0467 1 0.1548 1 1.0
Feature Patch RNN 0.0473 3 0.1564 2 2.5
Feature Token RNN 0.0471 2 0.1565 3 2.5
Temporal Token RNN 0.0480 4 0.1582 7 5.5
Temporal None MLP 0.0488 6 0.1580 6 6.0
Feature Patch RNN 0.0490 7 0.1577 5 6.0
Temporal Token MLP 0.0499 10 0.1575 4 7.0
Feature Token RNN 0.0483 5 0.1595 10 7.5
Temporal Freq MLP 0.0499 9 0.1582 9 9.0
Feature Patch Trans. 0.0493 8 0.1598 11 9.5
PEMS03_36_S Feature Patch RNN 0.0575 1 0.1678 1 1.0
Feature Patch Trans. 0.0586 2 0.1728 5 3.5
Temporal None MLP 0.0593 4 0.1715 4 4.0
Feature Patch RNN 0.0607 6 0.1690 2 4.0
Feature Token RNN 0.0598 5 0.1713 3 4.0
Temporal Token RNN 0.0587 3 0.1736 6 4.5
Temporal None MLP 0.0614 7 0.1741 7 7.0
Temporal Freq MLP 0.0634 9 0.1760 9 9.0
Feature Token RNN 0.0643 10 0.1758 8 9.0
Feature Patch Trans. 0.0619 8 0.1797 11 9.5
PEMS03_48_S Feature Patch RNN 0.0675 2 0.1771 1 1.5
Temporal None MLP 0.0702 4 0.1807 3 3.5
Feature Patch Trans. 0.0672 1 0.1858 7 4.0
Feature Patch RNN 0.0720 7 0.1778 2 4.5
Feature Token RNN 0.0714 5 0.1843 6 5.5
Temporal None MLP 0.0741 9 0.1827 4 6.5
Feature Patch Trans. 0.0701 3 0.1896 10 6.5
Feature Token RNN 0.0739 8 0.1840 5 6.5
Temporal Token RNN 0.0718 6 0.1875 8 7.0
Temporal Token RNN 0.0778 11 0.1894 9 10.0
Table 25: Top-10 configurations for the PEMS03 dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
PEMS04_12_S Temporal Freq MLP 0.0483 1 0.1638 2 1.5
Temporal Freq MLP 0.0488 3 0.1637 1 2.0
Feature Invert Trans. 0.0485 2 0.1646 3 2.5
Feature Token RNN 0.0495 4 0.1648 4 4.0
Temporal Invert MLP 0.0496 5 0.1656 5 5.0
Feature Patch Trans. 0.0496 6 0.1658 6 6.0
Feature Patch RNN 0.0503 7 0.1660 7 7.0
Temporal Invert MLP 0.0508 11 0.1670 8 9.5
Feature Invert RNN 0.0507 10 0.1675 9 9.5
Temporal None MLP 0.0506 9 0.1678 11 10.0
PEMS04_24_S Feature Patch Trans. 0.0607 1 0.1838 1 1.0
Feature Invert Trans. 0.0655 2 0.1838 2 2.0
Feature Token RNN 0.0687 4 0.1866 3 3.5
Feature Invert Trans. 0.0700 6 0.1916 7 6.5
Feature None Trans. 0.0703 9 0.1918 8 8.5
Feature Token RNN 0.0713 12 0.1914 6 9.0
Feature Invert RNN 0.0711 10 0.1926 10 10.0
Temporal Patch Trans. 0.0669 3 0.1952 18 10.5
Feature Patch RNN 0.0720 17 0.1898 4 10.5
Feature Patch Trans. 0.0719 16 0.1906 5 10.5
PEMS04_36_S Feature Patch RNN 0.0652 2 0.1867 1 1.5
Feature Invert Trans. 0.0645 1 0.1877 3 2.0
Feature Patch Trans. 0.0652 3 0.1874 2 2.5
Feature Patch RNN 0.0659 4 0.1885 4 4.0
Temporal Freq MLP 0.0667 5 0.1909 5 5.0
Feature Invert Trans. 0.0686 6 0.1936 6 6.0
Feature Patch Trans. 0.0687 7 0.1948 7 7.0
Temporal Invert MLP 0.0702 8 0.1957 9 8.5
Feature Token RNN 0.0711 10 0.1954 8 9.0
Temporal None MLP 0.0708 9 0.1958 10 9.5
PEMS04_48_S Feature Patch Trans. 0.0670 1 0.1900 1 1.0
Feature Patch RNN 0.0709 3 0.1939 2 2.5
Feature Patch RNN 0.0706 2 0.1949 3 2.5
Feature Patch Trans. 0.0721 4 0.1970 4 4.0
Feature Invert Trans. 0.0729 5 0.1978 5 5.0
Feature Token RNN 0.0750 6 0.2004 6 6.0
Feature Token RNN 0.0758 7 0.2007 7 7.0
Feature Invert Trans. 0.0763 8 0.2043 9 8.5
Temporal None MLP 0.0771 10 0.2027 8 9.0
Temporal Freq MLP 0.0766 9 0.2043 10 9.5
Table 26: Top-10 configurations for the PEMS04 dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
PEMS07_12_S Temporal Token RNN 0.0795 1 0.2047 1 1.0
Feature Patch Trans. 0.0823 2 0.2080 3 2.5
Feature Token RNN 0.0830 3 0.2079 2 2.5
Temporal Freq MLP 0.0834 4 0.2095 5 4.5
Temporal None MLP 0.0834 5 0.2096 6 5.5
Feature Invert RNN 0.0839 7 0.2093 4 5.5
Feature Patch Trans. 0.0835 6 0.2111 7 6.5
Feature None Trans. 0.0848 8 0.2114 9 8.5
Feature Invert Trans. 0.0857 10 0.2112 8 9.0
Feature None RNN 0.0856 9 0.2122 10 9.5
PEMS07_24_S Feature Patch RNN 0.1015 1 0.2292 1 1.0
Feature Patch RNN 0.1059 2 0.2331 2 2.0
Temporal Invert MLP 0.1085 4 0.2333 3 3.5
Feature Patch Trans. 0.1065 3 0.2347 4 3.5
Temporal Invert MLP 0.1112 6 0.2368 5 5.5
Temporal Token RNN 0.1087 5 0.2410 8 6.5
Temporal None MLP 0.1121 9 0.2395 6 7.5
Feature Invert RNN 0.1113 7 0.2418 9 8.0
Temporal Freq MLP 0.1136 10 0.2396 7 8.5
Temporal Token Trans. 0.1117 8 0.2420 10 9.0
PEMS07_36_S Feature Patch RNN 0.1241 2 0.2491 1 1.5
Temporal Token RNN 0.1204 1 0.2560 4 2.5
Feature Patch Trans. 0.1284 3 0.2554 3 3.0
Temporal Invert MLP 0.1321 6 0.2530 2 4.0
Temporal None MLP 0.1310 4 0.2575 6 5.0
Feature Patch RNN 0.1319 5 0.2570 5 5.0
Temporal Invert MLP 0.1346 7 0.2587 8 7.5
Feature Token RNN 0.1362 9 0.2583 7 8.0
Feature Token Trans. 0.1358 8 0.2625 11 9.5
Feature None Trans. 0.1366 11 0.2619 10 10.5
PEMS07_48_S Temporal None MLP 0.1408 2 0.2677 2 2.0
Feature Patch RNN 0.1429 3 0.2661 1 2.0
Temporal Token RNN 0.1349 1 0.2744 6 3.5
Temporal Invert MLP 0.1449 4 0.2716 4 4.0
Feature Patch RNN 0.1501 9 0.2682 3 6.0
Feature Patch Trans. 0.1466 6 0.2775 10 8.0
Feature Invert RNN 0.1462 5 0.2800 11 8.0
Feature None Trans. 0.1502 10 0.2748 8 9.0
Temporal Token Trans. 0.1490 8 0.2803 12 10.0
Feature Invert Trans. 0.1529 15 0.2741 5 10.0
Table 27: Top-10 configurations for the PEMS07 dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Setup IN SD Fusion Embed FF-Type MSE Rank MAE Rank Total
PEMS08_12_S Temporal Freq MLP 0.1575 1 0.2768 1 1.0
Temporal Token RNN 0.1597 2 0.2812 3 2.5
Feature Patch RNN 0.1608 3 0.2805 2 2.5
Feature None Trans. 0.1652 5 0.2847 6 5.5
Temporal Token Trans. 0.1659 7 0.2828 5 6.0
Feature Token RNN 0.1668 8 0.2822 4 6.0
Feature Patch Trans. 0.1653 6 0.2869 8 7.0
Feature Patch RNN 0.1671 9 0.2863 7 8.0
Temporal Patch MLP 0.1623 4 0.2902 15 9.5
Feature Invert Trans. 0.1696 11 0.2883 10 10.5
PEMS08_24_S Feature Patch RNN 0.1796 1 0.2985 1 1.0
Temporal Freq MLP 0.1802 2 0.3012 2 2.0
Feature Token RNN 0.1859 4 0.3029 4 4.0
Feature Patch Trans. 0.1851 3 0.3059 7 5.0
Temporal Token MLP 0.1866 6 0.3030 5 5.5
Temporal Invert MLP 0.1880 9 0.3028 3 6.0
Temporal Token Trans. 0.1880 8 0.3046 6 7.0
Feature Invert Trans. 0.1879 7 0.3070 11 9.0
Feature Patch RNN 0.1864 5 0.3078 13 9.0
Temporal None MLP 0.1911 11 0.3065 9 10.0
PEMS08_36_S Temporal Invert MLP 0.1980 2 0.3148 1 1.5
Feature Invert Trans. 0.1979 1 0.3163 3 2.0
Feature Token RNN 0.1994 3 0.3160 2 2.5
Feature Patch RNN 0.1994 4 0.3191 7 5.5
Feature Invert RNN 0.2018 7 0.3167 5 6.0
Temporal Token MLP 0.2036 10 0.3163 4 7.0
Feature Patch Trans. 0.2006 5 0.3198 9 7.0
Temporal None MLP 0.2028 9 0.3182 6 7.5
Feature Patch RNN 0.2012 6 0.3204 10 8.0
Temporal Freq MLP 0.2023 8 0.3206 11 9.5
PEMS08_48_S Temporal Token MLP 0.2059 3 0.3204 1 2.0
Temporal Invert MLP 0.2039 1 0.3236 5 3.0
Temporal None MLP 0.2055 2 0.3234 4 3.0
Feature Invert RNN 0.2073 5 0.3221 2 3.5
Feature Patch RNN 0.2078 6 0.3227 3 4.5
Feature Invert Trans. 0.2067 4 0.3241 6 5.0
Feature Token RNN 0.2099 8 0.3247 7 7.5
Feature Patch Trans. 0.2097 7 0.3283 8 7.5
Temporal Invert MLP 0.2129 10 0.3315 11 10.5
Feature None RNN 0.2155 12 0.3318 12 12.0
Table 28: Top-10 configurations for the PEMS08 dataset univariate forecasting. IN: Instance Norm, SD: Series Decomposition. ✓ indicates module used, ✗ indicates not used. Red/blue highlights indicate best and second-best performances.
Appendix D Limitation Discussion
In this work, we break down time-series models into five key design modules and benchmark a wide range of their combinations. Our framework identifies model configurations that can perform significantly better than existing standalone models. More importantly, our findings offer practitioners clear, understandable, and direct guidance. This ensures that the field advances with a deeper understanding and a more structured methodological approach, rather than through unsystematic exploration. However, we acknowledge that this work has the following limitations:
Not all time-series designs are covered. We recognize that TimeRecipe does not include every existing module design in time-series forecasting. The designs not covered generally fall into two categories. First, there are highly specialized designs, such as those in Crossformer [82]. These are specifically created for the Transformer architecture to capture dependencies across time and dimensions but are difficult to apply to other model architectures. Second, some model architectures are not yet widely adopted, such as models based on large language models (LLMs), for example, Time-LLM [26]. Such models often have high computational costs, and their effectiveness is still a subject of ongoing discussion [63].
Nevertheless, as the first-of-its-kind benchmark, TimeRecipe has covered a significant and representative portion of the design space for supervised time-series forecasting at the modular level.
The findings are primarily empirical. The insights from TimeRecipe are based on extensive empirical evaluations. While these experiments provide robust and thorough evaluations, and the toolkit recommends model architectures based on these empirical insights, a deeper theoretical analysis of why certain module designs are most effective under specific time-series forecasting scenarios is beyond the current scope of this benchmark study.
Nonetheless, our claims are supported by comprehensive benchmarking, over 10,000 experiments across diverse datasets, forecasting horizons, and task settings. The results are also statistically meaningful, with reported outcomes averaged over multiple random seeds for statistical robustness. Furthermore, statistical hypothesis testing validates the correlation findings.